### v2

# WRF-GS: Wireless Radiation Field Reconstruction with 3D Gaussian Splatting

Abstract
Wireless channel modeling plays a pivotal role in designing, analyzing, and optimizing wireless communication systems. Nevertheless, developing an effective channel modeling approach has been a longstanding challenge. This issue has been escalated due to the denser network deployment, larger antenna arrays, and wider bandwidth in 5G and beyond networks. To address this challenge, we put forth WRF-GS, a novel framework for channel modeling based on wireless radiation field (WRF) reconstruction using 3D Gaussian splatting. WRF-GS employs 3D Gaussian primitives and neural networks to capture the interactions between the environment and radio signals, enabling efficient WRF reconstruction and visualization of the propagation characteristics. The reconstructed WRF can then be used to synthesize the spatial spectrum for comprehensive wireless channel characterization. Notably, with a small number of measurements, WRF-GS can synthesize new spatial spectra within milliseconds for a given scene, thereby enabling latency-sensitive applications. Experimental results demonstrate that WRF-GS outperforms existing methods for spatial spectrum synthesis, such as ray tracing and other deep-learning approaches. Moreover, WRF-GS achieves superior performance in the channel state information prediction task, surpassing existing methods by a significant margin of more than 2.43 dB.
Index Terms:
Wireless channel modeling, wireless radiation field reconstruction, 3D Gaussian splatting, channel prediction.
I Introduction
Modern communications increasingly depend on wireless technologies that use electromagnetic (EM) waves for information exchange, driving advancements in mobile phones, automotive systems, and Internet-of-Thing devices [1]. At the core of these innovations lies wireless channel modeling, a long-standing problem in wireless communications. While the physics of EM wave propagation is described by Maxwell’s equations [2], solving these equations in reality is intricate due to the need for comprehensive knowledge of boundary conditions. This complexity has led to the development of various wireless channel modeling approaches, as shown in Fig. 1, which can be categorized into probabilistic modeling, deterministic modeling, and neural modeling.
The probabilistic models rely on statistical methods to predict channel characteristics, primarily estimating the received signal strength based on the distance between the transmitter (TX) and receiver (RX). These models are based on empirical formulas and use measurements to calibrate parameters for typical scenarios [3]. However, they often lack accuracy and struggle to provide detailed channel characteristics. For example, in multi-antenna systems, the probabilistic models fail to characterize the energy distribution of the received signal from all directions, i.e., to reconstruct the spatial spectrum of the received signal by estimating its angle of arrival (AoA). To overcome these drawbacks, the deterministic models use physical principles to predict channel characteristics under an approximate environmental model. For example, the ray tracing method generates propagation characteristics based on computer-aided design representations of the environment, including object boundaries and material reflection coefficients [4]. Therefore, these models can provide more comprehensive channel information than probabilistic models. However, their accuracy may be compromised, as they cannot capture the detailed physical characteristics of the environment.
Figure 1: Different types of wireless channel modeling.
In contrast, the neural models, adopting the data-driven principle, learn the complex interactions between the environment and radio signals directly from location-based data, rather than relying on predefined statistical or physical models. One of the promising advancements in neural models is the recent adoption of the neural radiance field (NeRF), a breakthrough in computer vision for view synthesis [5]. Motivated by the fact that light is a kind of EM wave, Refs. [6] and [7] proposed two NeRF-based frameworks, named NeRF2 and NeWRF, respectively, for wireless channel modeling based on implicit wireless radiation field (WRF) reconstruction. The reconstructed WRF can provide detailed and accurate channel characteristics and significantly enhance communication performance. However, these methods suffer from high computational complexity and slow synthesis (a.k.a. rendering) speeds. In practice, the synthesis operation corresponds to the channel characteristic prediction process and will significantly affect the round-trip time in communication systems, which is critical for latency-sensitive applications. For instance, cloud gaming and digital twins typically demand low latency, i.e., under 20 milliseconds for cloud gaming [8] and often within a few milliseconds for digital twins [9]. This renders the NeRF-based methods impractical.
To conquer this challenge, we propose WRF-GS, a hybrid framework for fast and accurate channel modeling based on explicit WRF reconstruction using neural networks and 3D Gaussian splatting (3D-GS). 3D-GS utilizes an analytical architecture to represent and render 3D scenes, enabling high-quality and real-time view synthesis [10]. Compared with NeRF, 3D-GS offers faster rendering speeds and lower computational complexity. By adapting 3D-GS from the optical domain to the radio-frequency (RF) domain, WRF-GS can predict the received signal at any location within milliseconds for a give scene after training with a small number of signal measurements. Fig. 2 illustrates the synthesized spatial spectra of four algorithms based on the reconstructed WRF in a laboratory environment when the RX is fixed at one position and the TX is located at four different positions (i.e., P1-P4). Evidently, WRF-GS has the best prediction accuracy compared with the baselines, closely matching the ground truth.
Figure 2: The synthesized spatial spectrum of different algorithms in a laboratory environment. The spatial spectrum represents the signal strength received from different directions composed of the azimuthal and elevation angles. (a) shows the 3D point clouds of the laboratory created by LiDAR, which are used for the ray-tracing algorithm. In addition, the TX can be located at any position, and the RX is equipped with a $4\times 4$ antenna array and fixed at a corner; (b) is a reconstructed WRF when the TX is located at P3 by running the WRF-GS framework using the 3D point clouds in (a) as initialization. The spatial spectrum is obtained by a series of nonlinear transformations, such as projecting the WRF onto the hemispherical plane, where the colors of the points represent different degrees of contribution to the spatial spectrum. (c) compares the synthesized spatial spectra of four algorithms when the TX is located at four different positions, i.e., P1-P4, as shown in (a). The ground truth is obtained using the antenna array.
Nevertheless, directly applying the optical 3D-GS technique to the RF domain poses several challenges. First, the optical 3D-GS only considers the amplitude (i.e., light strength) of the optical signal, while the received RF signal consists of both amplitude and phase components, as EM waves are more prone to reflection, diffraction, and scattering than lights. Second, the measurement of visible light is via a million-pixel camera, but an RX typically utilizes either a single antenna or an antenna array. Third, the rendering function in the RF domain differs from the optical 3D-GS due to the unique physics of EM wave propagation. To address these challenges, we propose WRF-GS, a novel framework that refines the optical 3D-GS technique [10] from three key aspects. Scenario Representation Network: We introduce two multi-layer perceptions (MLPs) to effectively capture the complex interactions between the environment and radio signals. The learned features are then embedded into 3D Gaussian points, which serve as virtual TXs in wireless propagation. Projection Model: We devise a projection model to transform the optical camera model into an RF antenna model using spherical and Mercator projections. This model projects the virtual TXs onto the RX perception plane. Electromagnetic Splatting: Based on the projected 3D Gaussian points, we employ a differentiable tile rasterizer to synthesize wireless channel characteristics. This operation enables rapid synthesis of new channel characteristics within milliseconds.
We conduct extensive experiments to evaluate the performance of the proposed WRF-GS. Unlike NeRF-based approaches that adopt an implicit representation, our WRF-GS explicitly represents the propagation characteristics, as shown in Fig. 2 (b). In addition, WRF-GS achieves higher reconstruction accuracy, higher sample efficiency, and faster rendering speed. Furthermore, we consider a downlink channel state information (CSI) prediction task in a multiple-input-multiple-output (MIMO) system as a case study to further demonstrate the effectiveness of WRF-GS in practical systems. Our results show that WRF-GS achieves better prediction performance of the downlink CSI than the baselines, outperforming NeRF2 by a significant margin of 2.43 dB.
In summary, this paper makes the following contributions: i) We propose WRF-GS, a hybrid framework tailored for wireless channel modeling based on WRF reconstruction using neural networks and 3D-GS; ii) We adapt the optical 3D-GS to the RF domain by designing the modules of scenario representation network, projection model, and electromagnetic splatting. The integration of the EM wave physics and 3D-GS enables a visible, low-cost, and accurate channel modeling approach; iii) We conduct extensive experiments and a case study to demonstrate the effectiveness of WRF-GS. Numerical results show that WRF-GS outperforms existing methods.
II Related Works
Wireless channel modeling involves measuring and characterizing the propagation of EM waves between TXs and RXs, taking into account various phenomena such as reflection, diffraction, scattering, and path loss [11, 12]. Sarkar et al. [3] surveyed various propagation models for mobile communications using statistical methods to describe the random nature of EM wave propagation. To provide more detailed channel characteristics, Oestges et al. [11] applied the ray tracing technique to simulate the paths of EM waves, which allows for precise predictions and accounts for large-scale features of the environment. These methods, however, fail to capture the materials and physical characteristics of the environment.
Environment-aware channel modeling has attracted much attention since future wireless networks will generate abundant location-specific channel data and are expected to possess more powerful data mining and AI capabilities [13, 14]. This motivated the adaptation of neural networks for more accurate channel modeling, i.e., learning the complex interactions between the environment and radio signals directly from location-based data [15]. To achieve this goal, a few recent works have been devoted to NeRF-based wireless channel modeling [16, 6, 7]. Orekondy et al. [16] proposed a neural surrogate to model wireless EM propagation effects in indoor environments based on NeRF using two synthetic datasets. Meanwhile, Zhao et al. [6] presented NeRF2 for spatial spectrum reconstruction in a real-world environment. By fixing a TX at an indoor location, Lu et al. [7] presented NeWRF, a NeRF-based wireless channel prediction framework with a sparse set of channel measurements. However, NeRF-based methods often suffer from high computational complexity and slower synthesis speed. In this paper, we present WRF-GS for environment-aware channel modeling with 3D-GS, reconciling the accuracy and computational complexity.
CSI prediction involves estimating future channel conditions based on current and historical data. Existing studies have been dedicated to predicting downlink channel states by observing uplink channels, as both are influenced by the same underlying physical environment and traverse identical paths [17, 18, 19]. Liu et al. [17] introduced FIRE, a variational auto-encoder (VAE) method, to transfer estimated CSI from the uplink to the downlink. In addition, Vasisht et al. [18] and Bakshi et al. [19] proposed R2F2 and OptML to predict the CSI using machine learning-based methods. In this paper, we apply WRF-GS to the CSI prediction task by transferring the estimated CSI from uplink to downlink.
III Preliminaries
In this section, we present preliminaries on wireless channel modeling and 3D Gaussian Splatting.
III-A Wireless Channel Modeling
A generic wireless communication system typically consists of a TX that generates and modulates an information signal, which is then transmitted through the wireless channel to an RX. The transmitted signal can be represented as a complex number $s=Ae^{j\varphi}$, where $A$ and $\varphi$ denote the amplitude and phase of the transmitted signal, respectively. In the wireless medium, the propagation of EM signals is characterized by key phenomena such as path loss and multipath propagation. For the basic free space path loss model, the received signal $y$ at the RX is given by
$y=Ae^{j\varphi}\cdot\Delta Ae^{j\Delta\varphi},$ (1)
where $\Delta A$ and $\Delta\varphi$ are the amplitude attenuation factor and phase rotation experienced by the signal, respectively. Due to various propagation effects, including reflection, scattering, refraction, and diffraction, the received signal is the sum of multiple copies of the transmitted signal. Consequently, it is more generally expressed as
$y=Ae^{j\varphi}\sum_{l=0}^{L-1}\Delta A_{l}e^{j\Delta\varphi_{l}},$ (2)
where $L$ is the total number of propagation paths, and $\Delta A_{l}$ and $\Delta\varphi_{l}$ represent the amplitude attenuation factor and phase rotation associated with the $l$-th path, respectively.
In multi-antenna systems, an antenna array is employed to characterize the energy distribution of the received signal $y$ from all directions, i.e., to estimate the AoA of the received signal. To illustrate this concept, we consider an antenna array equipped with $\sqrt{K}\times\sqrt{K}$ antennas, as depicted in Fig. 3. The spacing of two adjacent antennas is $D$, which is less than the wavelength $\lambda$. As shown in Fig. 3(b), the antenna array receives signals in a hemispherical plane to provide omnidirectional coverage and to efficiently capture signals from various directions. The direction of an RF source is characterized by the azimuthal angle $\alpha$ and elevation angle $\beta$, where $0^{\circ}\leq\alpha<360^{\circ}$ and $0^{\circ}\leq\beta<90^{\circ}$. The phase difference between the signals received by the $(m,n)$-th antenna pair is related to the distance of the two antennas from the source [20], i.e.,
$\Delta\theta_{m,n}$ $=\mathrm{mod}\left(\frac{2\pi\Delta d_{m,n}}{\lambda},2\pi\right)$
$=\mathrm{mod}\left(\frac{-2\pi r_{m,n}\cos\left(\alpha-\phi_{m,n}\right)\cos(\beta)}{\lambda},2\pi\right),$
where $\Delta d_{m,n}$ is shown in Fig. 3(a) and can be computed by
$\Delta d_{m,n}=|BA_{m,n}|-|BA_{0,0}|=-r_{m,n}\cos(\alpha-\phi_{m,n})\cos\beta,$
and $m,n=0,\ldots,\sqrt{K}-1$. In addition, $r_{m,n}=D\sqrt{m^{2}+n^{2}}$ and $\phi_{m,n}=\arctan 2(n,m)$ are the radius and angle of antenna $A_{m,n}$ at the polar coordinate system, respectively.
Figure 3: The AoA computation of the received signal with an antenna array. (a) There are $K$ antennas deployed in a grid, and the signal source is located at point B. (b) The spatial spectrum is generated from the antenna array. The illustration of spatial spectrum in Fig. 2(c) is obtained by projecting the spatial spectrum to the X-Y plane.
The AoA of the received signal is calculated by comparing the phases of signals received from multiple antennas. For AoA estimation, we use the antenna array to form a very narrow beam and steer it. When the beam is steered in the line-of-sight (LoS) direction, the array effectively blocks power from all other directions. This allows to reconstruct the spatial spectrum by steering the beam across a grid of azimuth and elevation angles and measuring the received signal power at each pointing direction. Specifically, we can form a matrix $\mathbf{P}$, where each element $\mathrm{P}(\alpha,\beta)$ represents the power of the received signal from the azimuthal angle $\alpha$ and elevation angle $\beta$, i.e.,
$\mathrm{P}(\alpha,\beta)=\left|\frac{1}{K}\sum_{m,n=0}^{\sqrt{K}-1,\sqrt{K}-1}e^{{j}\left(\Delta\hat{\theta}_{m,n}-\Delta\theta_{m,n}\right)}\right|^{2},\$ (3)
and
$\Delta\hat{\theta}_{m,n}=\hat{\theta}_{m,n}-\hat{\theta}_{0,0},$
where $\hat{\theta}_{m,n}$ denotes the measured phase of the RF signal received at the antenna $A_{m,n}$, and $\hat{\theta}_{0,0}$ is a constant term. Considering a one-degree angular resolution, the spatial spectrum can be reconstructed as a $360\times 90$ matrix:
[…]
In summary, the reconstructed spatial spectrum can be interpreted as a spatial power distribution function, which directly corresponds to the WRF emitting from the transmitting source and propagating through the environment. In other words, the spatial spectrum serves as a representation of the underlying WRF, providing valuable insights into wireless channel modeling. Therefore, this paper aims to reconstruct the WRF and synthesize the spatial spectrum using the 3D-GS technique, which is introduced in the next subsection.
III-B 3D Gaussian Splatting
3D-GS is a technique that utilizes a set of 3D Gaussian functions to model the geometry and appearance of a scene, offering a compact and efficient representation [10]. Unlike the computationally intensive and memory-heavy NeRF approach [5], 3D-GS provides a more lightweight and scalable solution for scene representation. Mathematically, 3D-GS provides a novel way to represent space using anisotropic Gaussian primitives, where each Gaussian is characterized by its own covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$ and center position vector $\boldsymbol{\mu}\in\mathbb{R}^{3}$. The 3D Gaussian distribution can be expressed as
$G(\boldsymbol{x})={e}^{-\frac{1}{2}(\boldsymbol{x}-\boldsymbol{\mu})^{T}\boldsymbol{\Sigma}^{-1}(\boldsymbol{x}-\boldsymbol{\mu})},$ (5)
where $\boldsymbol{x}=(x_{0},x_{1},x_{2})$ represents the three-dimensional spatial position of the Gaussian point, and $\boldsymbol{\Sigma}$ can be expressed in terms of a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ as
$\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}.$ (6)
The rendering process begins by transforming the 3D Gaussians $G(\boldsymbol{x})$ in space into 2D Gaussians $G^{\prime}(\boldsymbol{x^{\prime}})$ on the image plane [21]. These 2D Gaussians are then sorted based on their depth information. Each tile is processed independently and in parallel for the 2D Gaussians in its coverage. Utilizing the opacity ${o}_{i}$ and color attribute $c_{i}$ of each Gaussian, alpha blending is applied to the color values on each pixel:
$C=\displaystyle\sum_{i=1}^{N}{c}_{i}{\alpha}_{i}\displaystyle\prod_{j=1}^{i-1}(1-{\alpha}_{j})\ \mathrm{and}\ {\alpha}_{i}={o}_{i}{G}_{i}^{{}^{\prime}}(\varDelta\boldsymbol{p}_{i}),$ (7)
where $N$ is the number of 3D Gaussians near a given pixel, and $\varDelta\boldsymbol{p}_{i}=\boldsymbol{p}_{i}-\boldsymbol{p}_{s}$ is the difference vector between the projected center $\boldsymbol{p}_{i}$ of a Gaussian and the sampling pixel position $\boldsymbol{p}_{s}$. Once the color of each pixel is computed, the entire image is rendered. Subsequently, a comparison is made with the ground truth image to calculate the pixel-wise loss, which is then used to optimize the parameters of the model.
IV The WRF-GS Framework
In this section, we present the WRF-GS framework for wireless channel modeling based on WRF reconstruction using 3D-GS. This novel framework harnesses the power of 3D-GS and adapts it from the optical domain to the RF domain. The key insight is to convert the space particles into virtual TXs using a continuous volumetric scene function in 3D-GS.
IV-A Problem Description and Framework Overview
We consider a scenario comprising a TX, an RX equipped with an antenna array, and various obstacles such as floors, walls, and furniture. The RX is placed at a fixed position, while the TX is randomly positioned in the space. Following [6], we assume that the distribution of main obstacles in the space is fixed, and any temporary perturbations caused by moving small obstacles are mitigated through techniques like Kalman filtering to minimize their impact. As described in Section III-A, the received signal is the superposition of sub-signals from multiple propagation paths. Each sub-signal can be represented as an LoS signal from a virtual TX. Our objective is to reconstruct the spatial power spectrum received at the RX antenna array by synthesizing the distribution of these virtual TXs within the scenario.
Figure 4: An overview of the WRF-GS framework. The 3D point clouds, which can be randomly distributed or captured via LiDAR sensors, and the position of the TX are first passed into a scenario representation network. This network represents the virtual TXs in the scene using a set of 3D Gaussians, each of which carries environmental attenuation information and signal characteristics. To project these 3D Gaussian representations onto the perception plane of the RX antenna array, the Mercator projection is employed. Finally, the resulting spatial spectra are rendered using the electromagnetic splatting method.
This task is analogous to the optical 3D-GS technique that explicitly models the distribution of particles in a scene to enable arbitrary view synthesis [10]. This motivates us to employ 3D-GS to reconstruct the WRF and propose a novel framework named WRF-GS. Specifically, WRF-GS aims to model the virtual TX distribution to synthesize the received spatial spectrum by combining the EM wave physics and 3D-GS. Since 3D Gaussians possess attributes such as volume, color, and opacity, and serve as the fundamental building blocks in the representation space, they can be naturally translated into virtual TXs in the RF domain. During EM propagation, the overall outcome is perceived as the combined effect of the signal and attenuation, as described in Eqn. (2). In this context, the color attribute is replaced by the signal strength, while the opacity corresponds to the attenuation.
Nevertheless, the projection and splatting of WRF reconstruction in 3D-GS are distinct from the optical domain described in Section III-B. In optical 3D-GS, the goal is to project 3D Gaussian particles representing objects onto the camera’s 2D image plane. This involves applying a coordinate projection transformation that accounts for the camera’s position, orientation, and intrinsic parameters. In contrast, the RF domain requires mapping the virtual TXs represented by 3D Gaussians onto a perception plane of the RX antenna array. Since the RX antenna array receives signals in a hemispherical plane, as discussed in Section III-A, a new projection approach is required to map the 3D Gaussians onto a hemispherical plane, rather than a flat 2D plane. This operation is crucial for aligning the virtual TXs with the specific direction of the RX antenna array. Furthermore, in optical 3D-GS, the rendered image is generated by the splatted Gaussians, including volume, color, and opacity. While the 3D Gaussian attributes in the RF domain represent the signal strength and attenuation. This necessitates a new splatting process to blend the contributions of all the virtual TXs, accounting for their signal strengths and attenuation, to synthesize the received spatial power spectrum.
Our proposed WRF-GS effectively addresses the above challenges and enables the reconstruction of the spatial spectrum. The overall structure of WRF-GS is illustrated in Fig. 4. It takes the positions of physical TXs and 3D point clouds as input and outputs the spatial spectrum at the RX antenna array. WRF-GS consists of the following three core modules:
  • Scenario Representation Network: This module is responsible for representing the virtual TXs using 3D Gaussians. By processing the input data, the 3D Gaussians capture the signal strength and attenuation properties, which enables the synthesis of the WRF in the environment.
  • Projection Model: This model projects the virtual TXs represented by 3D Gaussians onto the perception plane of the RX antenna array. To account for the hemispherical nature of the antenna’s reception, we employ the Mercator projection to achieve this spatial projection.
  • Electromagnetic Splatting: Based on the properties of 3D Gaussians and the characteristics of EM wave propagation in space, we employ a hardware acceleration algorithm to efficiently process and combine the contributions of the virtual TXs onto the RX antenna array, ultimately synthesizing the received spatial spectrum.
In the following, we elaborate on these three modules.
IV-B Scenario Representation Network
The scenario representation network plays a crucial role in modeling the underlying relationships between the environment and radio signals by representing the virtual TXs as 3D Gaussians using two MLPs. The network architecture, depicted in Fig. 5, follows the DeepSDF structure [22].
To capture the attenuation, which solely depends on the spatial location, we initially feed the positions of the initial point clouds (equivalent to the center coordinates of the 3D Gaussians $G(\boldsymbol{x})$) into a network consisting of eight fully connected layers. Each layer comprises ReLU activations and 128 channels. The output of this network provides the attenuation $\delta(\boldsymbol{x})$ of each input 3D Gaussain $G(\boldsymbol{x})$ at the corresponding position $\boldsymbol{x}$ and a feature vector. Subsequently, we pass this feature vector, along with the TX position $P_{\text{TX}}$, through another two-layer fully connected network (with ReLU activations and 128 and 64 channels, respectively). The final output is the signal $S(\boldsymbol{x})$ that represents the 3D Gaussian signal associated with the virtual TX at position $\boldsymbol{x}$. Therefore, for each 3D Gaussian $G(\boldsymbol{x})$ representing a virtual TX, we obtain the corresponding signal $S(\boldsymbol{x})$ and attenuation $\delta(\boldsymbol{x})$ from the scenario representation network, i.e.,
${F}_{\Theta}:(G(\boldsymbol{x}),P_{\text{TX}})\Rightarrow\left(\delta(\boldsymbol{x}),S(\boldsymbol{x})\right),$ (8)
where $\Theta$ denotes the learnable neural network weights. Considering that signals and attenuation in space are typically expressed as complex numbers, as shown in Eqn. (1), the two outputs here are denoted by $\delta(\boldsymbol{x})=\Delta A(\boldsymbol{x})e^{j\Delta\psi(\boldsymbol{x})}$ and $S(\boldsymbol{x})=A(\boldsymbol{x})e^{j\psi(\boldsymbol{x})}$, respectively, where $A(\boldsymbol{x})$ represents the amplitude and $\psi(\boldsymbol{x})$ represents the phase.
This network is specifically designed to transform the 3D point clouds and TX position information into the signal and attenuation properties of 3D Gaussians, respectively. Unlike DeepSDF, which is typically applied to optics and assumes a fixed TX (light source) position, our network is designed to handle mobile TXs. Moreover, the outputs are all complex-valued signals that incorporate additional phase information. This unique feature enables our model to successfully represent both the signal characteristics of virtual TXs in space and the attenuation information of the surrounding environment.
Figure 5: Architecture of the neural model. The model comprises two MLPs. The first MLP processes the input 3D point clouds to capture the spatial attenuation information of the environment, which is independent of the specific TX position. This MLP outputs the attenuation and a feature vector for each spatial location. The second MLP then combines the spatial feature vector with the TX position to capture the signal characteristics.
IV-C Projection Model
The projection model maps the virtual TXs represented by 3D Gaussians onto the perception plane of the RX antenna array. There are several key differences compared with the camera model in optical 3D-GS. Unlike pinhole or fisheye cameras, the antenna in our WRF-GS receives signals from a hemispherical direction. In addition, we do not need to know the distortion caused by coordinate transformation. This is because we only care about the signal at the corresponding angular resolution and ignore the signal between angles. Moreover, as shown in Eqn. (4), the arrangement of pixels in the final spatial spectrum differs from conventional photographs that adhere to human visual perception.
To project a 3D point $\boldsymbol{t}={[{t}_{x},{t}_{y},{t}_{z}]}^{T}$ onto the perception plane of the RX antenna array at $\boldsymbol{p}={[{p}_{x},{p}_{y}]}^{T}$, we employ the Mercator projection, a widely used orthographic cylindrical projection in cartography. This allows us to map the latitude and longitude grid of a sphere onto a cylindrical surface, which can then be unrolled onto a flat plane. We define a right-handed Cartesian coordinate at the receiving end of the antenna, with the antenna array oriented in the positive Z-direction and the signal being received in a spherical plane, as shown in Fig. LABEL:Proj(a). Using longitude $\Omega_{\text{lon}}$ and latitude $\Omega_{\text{lat}}$, we can concisely and accurately represent positional relationships in space, which is expressed using inverse trigonometric functions as:
$\begin{bmatrix}\Omega_{\text{lon}}\\ \Omega_{\text{lat}}\end{bmatrix}=\begin{bmatrix}\arctan 2({t}_{y}/{t}_{x})\\ \arcsin({t}_{z}/{t}_{r})\end{bmatrix},$ (9)
where ${t}_{r}=\sqrt{{t}_{x}^{2}+{t}_{y}^{2}+{t}_{z}^{2}}$ denotes the distance from the center coordinate of the 3D Gaussian in space to the center of the antenna coordinate, $\arctan 2(\cdot)$ is the 4-quadrant inverse tangent. Since we only consider signals from the upper hemisphere, we have $-\pi\leq\Omega_{\text{lon}}<\pi$ and $20\leq\Omega_{\text{lat}}<\pi/$ 2, and the lower hemisphere is ignored in the following coordinate transformations. Subsequently, we convert the latitude-longitude coordinates (see Fig. LABEL:Proj(b)) into uniform coordinates (see Fig. LABEL:Proj(c)) as
$\begin{bmatrix}{s}_{x}\\ {s}_{y}\end{bmatrix}=\begin{bmatrix}\Omega_{\text{lon}}/\pi\\ 2\Omega_{\text{lat}}/\pi\end{bmatrix},$ (10)
such that $-1\leq{s}_{x}<1$ and $10\leq{s}_{y}<$ 1. The final step of the projection is to transform the uniform coordinates into pixel coordinates of arbitrary resolution (see Fig. LABEL:Proj(d)) as
$\begin{bmatrix}{p}_{x}\\ {p}_{y}\end{bmatrix}=\begin{bmatrix}({s}_{x}+1)\times W/2\\ {s}_{y}\times H\end{bmatrix},$ (11)
where $W$ and $H$ represent the number of pixels contained in the width and height of the image, respectively. Considering a one-degree angular resolution, we set $W=360$ and $H=90$.
Through the above coordinate transformations, we maintain the relationship between the signal angles, ensuring that each pixel value in the spatial spectrum corresponds to the AoA of the signal. This allows us to calculate the signals arriving in each direction at the RX antenna array in the next subsection.
IV-D Electromagnetic Splatting
After the coordinate projection transformation, the 3D Gaussians representing the virtual TXs are mapped onto the 2D plane. Note that each 2D Gaussian may cover multiple pixels. In other words, the value of each pixel is the result of the combined contributions from multiple 2D Gaussians. Therefore, the role of electromagnetic splatting module is to efficiently compute the value of each pixel in the 2D plane, which is then utilized to reconstruct the spatial spectrum.
As illustrated in Fig. 9, electromagnetic splatting incorporates the EM wave propagation into the differentiable rasterisation algorithm [10]. To facilitate fast parallel computing, we group adjacent pixels into non-overlapping “tiles” that can be computed independently, as depicted in Fig. 9(a). For each 2D Gaussian, we first determine which tiles it covers, duplicate the Gaussian, and assign the copies to the corresponding tiles. These 2D Gaussians are then sorted by depth order within each tile, as shown in Fig. 9(b). In the scenario representation network, we obtain the signal $S(\boldsymbol{x}_{i})$ and signal attenuation $\delta(\boldsymbol{x}_{i})$ of each 3D Gaussian $G(\boldsymbol{x}_{i})$ at an arbitrary position $\boldsymbol{x}_{i}$. Assuming there are a total of $N$ 3D Gaussians contributing to a particular angle, we can express the effect of the $i^{th}$ Gaussian $G(\boldsymbol{x}_{i})$ on the RX by combining Eqn. (2) and Eqn. (7) as
$S_{i}(\boldsymbol{x})=\left(\prod_{j=0}^{i-1}\delta(\boldsymbol{x}_{j})\right)S(\boldsymbol{x}_{i}).$ (12)
Consequently, for all 3D Gaussians associated with an angle (pixel) $k$ , the combined effect is expressed as
$R_{k}=\sum_{i=1}^{N}S_{i}(\boldsymbol{x}).$ (13)
To facilitate a more intuitive understanding of this process, we provide an illustrative example in Fig. 9(c). For the pixel value $R_{1}$ in the upper-left corner, only the signal $S(\boldsymbol{x}_{1})$ from $G(\boldsymbol{x}_{1})$ contributes, without any attenuation. For the pixel value $R_{2}$ in the upper-right corner, the contributions from $G(\boldsymbol{x}_{1})$, $G(\boldsymbol{x}_{2})$, and $G(\boldsymbol{x}_{3})$ must be considered. The signal from $G(\boldsymbol{x}_{1})$ has no attenuation, but the signals from $G(\boldsymbol{x}_{2})$ and $G(\boldsymbol{x}_{3})$ are attenuated by $\delta(\boldsymbol{x}_{1})$ and $\delta(\boldsymbol{x}_{1})\delta(\boldsymbol{x}_{2})$, respectively, due to the depth ordering. The computation for the remaining pixel values $R_{k}$ follows a similar principle.
Figure 9: An illustration of electromagnetic splatting. (a) The 3D Gaussians are splatted onto a 2D plane, and the plane is divided into non-overlapping tiles to enable parallel computation. (b) Gaussians that cover multiple tiles are duplicated and assigned to the corresponding tiles, then sorted by depth order within each tile. (c) The value of each pixel can be calculated independently and in parallel using Eqn. (12) and Eqn. (13).
In simple terms, we consider each projected Gaussian as a virtual TX that can affect multiple angles, while signals from multiple virtual TXs are received at each angle. Depending on the distance of these signals to the RX, they will be attenuated to different degrees, ultimately resulting in the received signal on each angle. As explained in Section III-A, the reconstructed spatial spectrum can be represented as a spatial power distribution function, where the value is proportional to the square of the amplitude of the received signal. Therefore, by deriving the received signal on each angle from Eqn. (12) and Eqn. (13), we obtain the spatial spectrum described in Eqn. (4), which serves as a representation of the underlying WRF.
IV-E Summary
The WRF-GS framework integrates physical and deep learning models to enable comprehensive environment-aware channel modeling and high-fidelity spatial spectrum synthesis. At the core of this approach lies the establishment of a connection between virtual TXs and 3D Gaussians. The scenario representation network models the spatial distribution of signal propagation and attenuation using 3D Gaussians based on input data. To transform these 3D spatial representations onto a 2D plane for further processing, we employ the Mercator map projection, ensuring accurate spatial mapping. Finally, the electromagnetic splatting module employs a parallel and efficient algorithm to compute pixel values in each direction, thereby achieving fast spatial spectrum synthesis.
V WRF-GS Implementation and Evaluation
In this section, we implement and evaluate WRF-GS in a laboratory environment using a real-world dataset in [6].
V-A Implementation
WRF-GS is implemented in a Python environment using modified custom CUDA kernels for rasterization [10] and is trained on an NVIDIA GeForce RTX 3090 GPU. Detailed settings and implementation are provided below.
Position Initialization: We use the LiDAR 3D point clouds [^1] Note that WRF-GS also works well by randomly initializing the point clouds within a specified range of space. shown in Fig. 2(a) to initialize the positions of 3D Gaussian points. The reason is that the most significant signal attenuation usually occurs on the object’s surface [7], which corresponds to the 3D point clouds. This initialization enables a fast convergence of WRF-GS. Besides, we adopt an adaptive density control strategy [10] to learn the position distribution of the 3D Gaussians during the training process.
Position Encoding: To improve the spatial resolution, we introduce an efficient position encoding method [5], i.e.,
$\gamma(\boldsymbol{t})=(\sin(\pi\boldsymbol{t}),\cos(\pi\boldsymbol{t}),\ldots,\sin(2^{L}\pi\boldsymbol{t}),\cos(2^{L}\pi\boldsymbol{t})),$ (14)
where $\boldsymbol{t}$ represents the 3D coordinates of the TX and $L$ is the order of the position encoding. For the input position encoding of $P_{\text{TX}}$ and $G(\boldsymbol{x})$ in the MLPs, we set $L=9$.
Optimization Details: In the scenario representation network, we need to train two MLPs and learn the 3D Gaussian representations using the training dataset. We adopt the Adam optimizer for network training and 3D Gaussian representation. The loss function is defined as the difference between the synthesized spatial spectrum $I_{\text{pred}}$ and the ground truth $I_{\text{gt}}$, i.e.,
$\mathcal{L}=(1-\eta)|I_{\text{gt}}-I_{\text{pred}}|+\eta(1-\xi(I_{\text{gt}},I_{\text{pred}})),$ (15)
where the $\xi(I_{\text{gt}},I_{\text{pred}})$ is the structural similarity index measure (SSIM) function [23], which is used to measure the similarity between two images. In addition, $\eta=0.2$ is a weighting factor.
V-B Experiment Design
Dataset. We use an open-source dataset provided in NeRF2 [6] to evaluate the WRF-GS framework. This dataset was obtained in a laboratory environment, which consists of multiple tables, shelves, and several small rooms. The 3D LiDAR point clouds of the environment are shown in Fig. 2(a). The RX is equipped with a $4\times 4$ antenna array and operated at 915MHz, while the TX is an RFID tag that sends RN16 messages repeatedly. To collect the training dataset, the position of the RX is fixed and the position of the TX is systematically varied throughout the laboratory. Each sample in the dataset consists of the TX position and the corresponding spatial spectrum measured at the RX (i.e., Eqn (4)).
Benchmarks. We compare WRF-GS with several baselines:
  • Ray Tracing[2]: This approach emits rays from the RX through every pixel in the laboratory using the 3D LiDAR point clouds. These rays interact with objects to simulate the propagation of signal in the real world.
  • Variational Autoencoder (VAE) [24]: VAE is commonly used to generate new data samples by mapping observed data to a continuous latent space [25]. Here, we use VAE to predict the spatial spectrum at arbitrary TX positions.
  • Deep Convolutional Generative Adversarial Network (DCGAN) [26]: DCGAN comprises a generator and a discriminator that are trained in an adversarial manner. The generator learns to generate realistic samples to deceive the discriminator, while the discriminator aims to distinguish the real samples from the generated ones.
  • NeRF2 [6]: NeRF2 is a method for spatial spectrum synthesis using the NeRF technique. After training, NeRF2 can predict the spatial spectrum at arbitrary TX positions.
V-C Evaluation
We first evaluate the accuracy of the spatial spectrum synthesis generated by the WRF-GS, NeRF2, VAE, and ray tracing methods, as shown in Fig. 2(c). The visualization is obtained by transforming the matrix form in Eqn. (4) into a polar coordinate representation. The spatial spectrum of WRF-GS for the case when the TX is located at position P3 is a projection of the constructed WRF visualized in Fig. 2(b). The results in Fig. 2(c) show that WRF-GS achieves the best prediction accuracy, closely matching the ground truth.
Fig. 10 compares the SSIM [23] values of the synthesized spatial spectra generated by different methods. SSIM values range from 0 to 1, with higher values indicating greater similarity between the synthesized and groud truth samples. The cumulative distribution function (CDF) of SSIM illustrates the disparities between the methods and the ground truth. From Fig. 10, we see that WRF-GS outperforms the baselines. Specifically, the median SSIM values for WRF-GS, NeRF2, VAE, DCGAN, and ray tracing are 0.82, 0.78, 0.70, 0.56, and 0.38, respectively. The slightly inferior performance of NeRF2 is due to the challenge in enumerating the virtual TX locations along the propagation paths for integration or rendering. In contrast, the ray tracing method achieves the worst performance, as it heavily relies on accurate environmental material information.
Figure 10: The SSIM values of the synthesized spatial spectra generated by the WRF-GS, NeRF2, VAE, DCGAN, and ray tracing methods.
Figure 11: The median SSIM values of the WRF-GS and NeRF2 methods under different training dataset sizes.
Fig. 11 presents the median SSIM values of the WRF-GS and NeRF2 methods across varying sizes of the training dataset. We maintain a consistent test dataset and randomly subsample the original training dataset to create new training datasets ranging from $200$ to $4,800$ samples. The results demonstrate that the WRF-GS framework consistently outperforms the NeRF2 method for all training dataset sizes, improving the median SSIM by $4\%$ to $8\%$. More importantly, the performance gap increases as the training dataset size decreases. This suggests that WRF-GS is well-suited for practical applications in channel modeling, as it requires less training data and, consequently, less time and effort from experienced engineers.
Figure 12: The rendering time required to synthesize a spatial spectrum using different methods under the same settings.
Fig. 12 illustrates the rendering time required to synthesize a spatial spectrum using different methods under the same settings. The results show that, for processing a single sample, WRF-GS takes $0.005$s, while NeRF2, VAE, DCGAN, and Ray Tracing takes $0.2$s, $0.05$s, $0.09$s, and $7$s, respectively. The processing time of WRF-GS is usually one to three orders of magnitude less than other methods. This superior computational efficiency makes WRF-GS well-suited for latency-sensitive applications.
VI Case Study: CSI Prediction
In this section, we demonstrate the effectiveness of WRF-GS on the downlink CSI prediction task in MIMO systems. Typically, the downlink and uplink transmission are operated on different frequency bands in frequency domain duplex systems, where the uplink-downlink reciprocity does not hold. The base station (BS) can obtain the downlink CSI of each antenna from the uplink channel feedback by sending pilot sequences to the end devices. However, the overhead of such feedback is linearly proportional to the number of antennas and devices, making it impossible in massive MIMO systems.
Motivated by the fact that the uplink and downlink transmissions experience the same physical environment, it is reasonable to use the uplink CSI measured at the BS to infer the downlink CSI. To achieve this goal, we employ WRF-GS to predict the downlink CSI based on the uplink CSI. As pointed out in [27], each CSI is unique and highly correlated with the physical environment. Therefore, there is a mapping between the device’s location and its uplink CSI. Accounting for this, Eqn. (8) can be rewritten as
${F}_{\Theta}:(G(\boldsymbol{x}),I_{u}(\boldsymbol{x}))\Rightarrow\left(\delta(\boldsymbol{x}),S(\boldsymbol{x})\right),$ (16)
where $I_{u}(\boldsymbol{x})$ is the uplink CSI at location $\boldsymbol{x}$. Thus, we can retrain a WRF-GS to perform the downlink CSI prediction task. In this case, the input is changed from the TX’s location to the uplink CSI, and the output is changed from the spatial spectrum to the downlink CSI.
VI-A Experiment Setup
We adopt the public Argos channel dataset [28] to train the MLPs and 3D Gaussian representations in WRF-GS for the downlink CSI prediction task. This dataset is collected in a real environment and the CSI is measured in different environments where a BS is equipped with $104$ antennas and serves multiple users. Each CSI measurement contains $52$ subcarriers. Similar to previous works [6, 17], we regard the first $26$ subcarriers as the uplink channel and the remaining $26$ subcarriers as the downlink channel. Apart from NeRF2 [6], we also compare WRF-GS with the following three approaches:
  • FIRE [17]: FIRE is a deep learning method based on the VAE architecture and is used to predict the downlink CSI by learning the latent distribution of the uplink CSI.
  • R2F2 [18]: Given the uplink CSI, R2F2 can obtain each path’s parameters and the number of paths by solving an optimization problem to estimate the downlink CSI.
  • OptML [19]: OptML is a deep learning method based on the R2F2 framework and can be used to predict the underlying path information of the downlink channel.
VI-B Channel Estimation Accuracy
Fig. 13 depicts the prediction accuracy of the WRF-GS framework for the downlink CSI prediction task. The blue solid line represents the input ground truth uplink (GT Uplink) CSI, the orange solid line represents the ground truth downlink (GT Downlink) CSI, and the black dashed line is the predicted (Pred Downlink) CSI output by WRF-GS. It is seen that the predicted result is very close to the ground truth.
Figure 13: The comparison of the ground truth CSI and the predicted CSI using WRF-GS in terms of the channel amplitude and phase via different sub-carries.
To quantify the prediction accuracy, we use the SNR metric:
$\text{SNR}=-10\log_{10}\left({\left\|\mathrm{}C_{\text{pred}}-C_{\text{GT}}\right\|^{2}}/{\left\|C_{\text{GT}}\right\|^{2}}\right),$ (17)
where $C_{\text{pred}}$ is predicted downlink channel and $C_{\text{GT}}$ is ground truth downlink channel. The higher the SNR, the better the prediction accuracy. Fig. 14 shows the SNRs of five CSI prediction algorithms. Notably, WRF-GS achieves a median SNR of $22.98$ dB ($10$th percentile: $20.82$ dB, $90$th percentile: $24.33$ dB). The median SNRs of NeRF2, FIRE, R2F2, and OptML are $20.55$ dB, $15.29$ dB, $8.57$ dB, and $8.47$ dB, respectively. This demonstrates that WRF-GS improves the CSI prediction accuracy by efficiently modeling wireless channels.
Figure 14: The prediction accuracy of different methods in the downlink CSI prediction task. The figure shows the maximum, minimum, median, upper and lower quartile ranges of the SNR distribution for different methods, as well as the outliers. The higher the SNR, the better the prediction accuracy.
VII Conclusion
This paper introduced WRF-GS, a novel framework that combines 3D-GS and neural networks to enable fast and accurate wireless channel modeling. To adapt the 3D-GS technique from the optical domain to the RF domain, we designed three key modules for WRF-GS: a scenario representation network, a projection model, and an electromagnetic splatting technique. These modules allow WRF-GS to accurately capture the underlying physics of EM wave propagation. Remarkably, WRF-GS can synthesize new channel characteristics within milliseconds using a small number of training samples. Experimental results showed that WRF-GS outperforms existing methods in spatial spectrum reconstruction and excels in the CSI prediction task, achieving an improvement of over 2.43 dB. We envision that WRF-GS will enable various applications in future networks by leveraging its high-fidelity visualization, low sample complexity, and rapid rendering speed. The authors have provided public access to their code at https://github.com/wenchaozheng/WRF-GS.

### v4

@@
-WRF-GS: Wireless Radiation Field Reconstruction with 3D Gaussian Splatting
+Neural Representation for Wireless Radiation Field Reconstruction: A 3D Gaussian Splatting Approach
 Abstract
 Wireless channel modeling plays a pivotal role in designing, analyzing, and optimizing wireless communication systems.
-Nevertheless, developing an effective channel modeling approach has been a longstanding challenge.
-This issue has been escalated due to the denser network deployment, larger antenna arrays, and wider bandwidth in 5G and beyond networks.
-To address this challenge, we put forth WRF-GS, a novel framework for channel modeling based on wireless radiation field (WRF) reconstruction using 3D Gaussian splatting.
+Nevertheless, developing an effective channel modeling approach has been a long-standing challenge.
+This issue has been escalated due to denser network deployment, larger antenna arrays, and broader bandwidth in next-generation networks.
+To address this challenge, we put forth WRF-GS, a novel framework for channel modeling based on wireless radiation field (WRF) reconstruction using 3D Gaussian splatting (3D-GS).
 WRF-GS employs 3D Gaussian primitives and neural networks to capture the interactions between the environment and radio signals, enabling efficient WRF reconstruction and visualization of the propagation characteristics.
 The reconstructed WRF can then be used to synthesize the spatial spectrum for comprehensive wireless channel characterization.
-Notably, with a small number of measurements, WRF-GS can synthesize new spatial spectra within milliseconds for a given scene, thereby enabling latency-sensitive applications.
-Experimental results demonstrate that WRF-GS outperforms existing methods for spatial spectrum synthesis, such as ray tracing and other deep-learning approaches.
-Moreover, WRF-GS achieves superior performance in the channel state information prediction task, surpassing existing methods by a significant margin of more than 2.43 dB.
+While WRF-GS demonstrates remarkable effectiveness, it faces limitations in capturing high-frequency signal variations caused by complex multipath effects.
+To overcome these limitations, we propose WRF-GS+, an enhanced framework that integrates electromagnetic wave physics into the neural network design.
+WRF-GS+ leverages deformable 3D Gaussians to model both static and dynamic components of the WRF, significantly improving its ability to characterize signal variations.
+In addition, WRF-GS+ enhances the splatting process by simplifying the 3D-GS modeling process and improving computational efficiency.
+Experimental results demonstrate that both WRF-GS and WRF-GS+ outperform baselines for spatial spectrum synthesis, including ray tracing and other deep-learning approaches.
+Notably, WRF-GS+ achieves state-of-the-art performance in the received signal strength indication (RSSI) and channel state information (CSI) prediction tasks, surpassing existing methods by more than 0.7 dB and 3.36 dB, respectively.
+The code is available at https://github.com/wenchaozheng/WRF-GSplus.
 Index Terms:
@@
 I Introduction
-Modern communications increasingly depend on wireless technologies that use electromagnetic (EM) waves for information exchange, driving advancements in mobile phones, automotive systems, and Internet-of-Thing devices [1].
-At the core of these innovations lies wireless channel modeling, a long-standing problem in wireless communications.
-While the physics of EM wave propagation is described by Maxwell’s equations [2], solving these equations in reality is intricate due to the need for comprehensive knowledge of boundary conditions.
+Modern communications increasingly depend on wireless technologies that use electromagnetic waves (EM) for information exchange, driving advances in mobile phones, automotive systems, and the Internet of Things [2].
+At the core of these advances lies wireless channel modeling, a long-standing problem for modeling the interactions between the environment and radio signals in wireless communications.
+While Maxwell’s equations model such interactions [3], solving these equations in reality is intricate due to the need for comprehensive knowledge of boundary conditions.
 This complexity has led to the development of various wireless channel modeling approaches, as shown in Fig.
-1, which can be categorized into probabilistic modeling, deterministic modeling, and neural modeling.
+1, which are categorized into probabilistic modeling, deterministic modeling, and neural modeling.
 The probabilistic models rely on statistical methods to predict channel characteristics, primarily estimating the received signal strength based on the distance between the transmitter (TX) and receiver (RX).
@@
 Figure 1: Different types of wireless channel modeling.
-In contrast, the neural models, adopting the data-driven principle, learn the complex interactions between the environment and radio signals directly from location-based data, rather than relying on predefined statistical or physical models.
+In contrast, the neural models, adopting the data-driven principle, learn the complex interactions between the environment and radio signals directly from location-based data.
 One of the promising advancements in neural models is the recent adoption of the neural radiance field (NeRF), a breakthrough in computer vision for view synthesis [6].
@@
 This renders the NeRF-based methods impractical.
-To conquer this challenge, we propose WRF-GS, a hybrid framework for fast and accurate channel modeling based on explicit WRF reconstruction using neural networks and 3D Gaussian splatting (3D-GS).
-3D-GS utilizes an analytical architecture to represent and render 3D scenes, enabling high-quality and real-time view synthesis [10].
-Compared with NeRF, 3D-GS offers faster rendering speeds and lower computational complexity.
-By adapting 3D-GS from the optical domain to the radio-frequency (RF) domain, WRF-GS can predict the received signal at any location within milliseconds for a give scene after training with a small number of signal measurements.
+Figure 2: The synthesized spatial spectra of five algorithms in a laboratory environment.
+The spatial spectrum represents the signal strength received from different directions composed of the azimuthal and elevation angles.
+(a) shows the 3D point clouds of the laboratory created by LiDAR, which are used for the ray tracing algorithm.
+In this setting, the TX can be located at any position, while the RX is equipped with a $4\times 4$ antenna array and fixed at a corner.
+(b) compares the synthesized spatial spectra of five algorithms when the TX is located at four different positions, i.e., P1-P4, as shown in (a).
+The ground truth is obtained using the antenna array.
+Recently, 3D Gaussian splatting (3D-GS) has emerged as a promising technique for efficient and compact scene reconstruction using a set of 3D Gaussian functions [11].
+Compared with NeRF, 3D-GS offers significantly faster rendering speeds and lower computational complexity.
+These advantages motivate us to adapt the 3D-GS technique for reconstructing the WRF.
+Nevertheless, directly applying the optical 3D-GS to the RF domain poses several significant challenges.
+First, the optical 3D-GS only considers the amplitude of optical signals (i.e., light intensity).
+In contrast, the received RF signals consist of both amplitude and phase components, as EM waves are more prone to reflection, diffraction, and scattering than visible light.
+Second, visible light is typically measured using million-pixel cameras, while an RX employs either a single antenna or an antenna array.
+Third, the rendering function in the RF domain differs from that in optical 3D-GS due to the unique physics of EM wave propagation.
+To address these challenges, we propose WRF-GS, a novel framework for fast and accurate channel modeling based on explicit WRF reconstruction using neural networks and 3D-GS.
+Specifically, we refine the optical 3D-GS technique from three key aspects to adapt it to the RF domain.
+First, we design a scenario representation network that comprises two multilayer perceptions (MLPs) to effectively capture the complex interactions between the environment and radio signals.
+The learned features are then embedded into 3D Gaussian points, which serve as virtual TXs to model wireless signal propagation.
+This approach enables the network to represent the spatial distribution of signals.
+Second, we devise a projection model to adapt the optical camera model to an RF antenna model using spherical and Mercator projections.
+This model projects the virtual TXs onto the RX perception plane, enabling accurate spatial mapping of signal propagation.
+Third, we introduce an electromagnetic splatting process using a differentiable tile rasterizer to synthesize wireless channel characteristics based on the projected 3D Gaussian points.
+This process, combined with the projection model, enables rapid synthesis of new channel characteristics within milliseconds.
+However, WRF-GS faces limitations in accurately modeling high-frequency signal variations and also suffers from parameter redundancy due to its coarse neural network design.
+To overcome these limitations, we further put forth WRF-GS+, an enhanced version of WRF-GS that integrates electromagnetic wave physics into the neural network design and eliminates the need for additional parameters.
+Specifically, we adopt the deformable 3D Gaussians in the scenario representation network to model both static and dynamic components of the WRF, significantly improving the ability to capture the complex interactions between the environment and radio signals.
+In addition, we exploit the inherent properties of 3D Gaussian primitives directly and leverage the $\alpha$-blending technique to synthesize wireless channel characteristics based on the projected 3D Gaussian points.
+These enhancements enable WRF-GS+ to characterize the distribution of the WRF in a given environment more effectively and with greater computational efficiency.
 Fig.
-2 illustrates the synthesized spatial spectra of four algorithms based on the reconstructed WRF in a laboratory environment when the RX is fixed at one position and the TX is located at four different positions (i.e., P1-P4).
-Evidently, WRF-GS has the best prediction accuracy compared with the baselines, closely matching the ground truth.
-Figure 2: The synthesized spatial spectrum of different algorithms in a laboratory environment.
-The spatial spectrum represents the signal strength received from different directions composed of the azimuthal and elevation angles.
-(a) shows the 3D point clouds of the laboratory created by LiDAR, which are used for the ray-tracing algorithm.
-In addition, the TX can be located at any position, and the RX is equipped with a $4\times 4$ antenna array and fixed at a corner; (b) is a reconstructed WRF when the TX is located at P3 by running the WRF-GS framework using the 3D point clouds in (a) as initialization.
-The spatial spectrum is obtained by a series of nonlinear transformations, such as projecting the WRF onto the hemispherical plane, where the colors of the points represent different degrees of contribution to the spatial spectrum.
-(c) compares the synthesized spatial spectra of four algorithms when the TX is located at four different positions, i.e., P1-P4, as shown in (a).
-The ground truth is obtained using the antenna array.
-Nevertheless, directly applying the optical 3D-GS technique to the RF domain poses several challenges.
-First, the optical 3D-GS only considers the amplitude (i.e., light strength) of the optical signal, while the received RF signal consists of both amplitude and phase components, as EM waves are more prone to reflection, diffraction, and scattering than lights.
-Second, the measurement of visible light is via a million-pixel camera, but an RX typically utilizes either a single antenna or an antenna array.
-Third, the rendering function in the RF domain differs from the optical 3D-GS due to the unique physics of EM wave propagation.
-To address these challenges, we propose WRF-GS, a novel framework that refines the optical 3D-GS technique [10] from three key aspects.
-Scenario Representation Network: We introduce two multi-layer perceptions (MLPs) to effectively capture the complex interactions between the environment and radio signals.
-The learned features are then embedded into 3D Gaussian points, which serve as virtual TXs in wireless propagation.
-Projection Model: We devise a projection model to transform the optical camera model into an RF antenna model using spherical and Mercator projections.
-This model projects the virtual TXs onto the RX perception plane.
-Electromagnetic Splatting: Based on the projected 3D Gaussian points, we employ a differentiable tile rasterizer to synthesize wireless channel characteristics.
-This operation enables rapid synthesis of new channel characteristics within milliseconds.
-We conduct extensive experiments to evaluate the performance of the proposed WRF-GS.
-Unlike NeRF-based approaches that adopt an implicit representation, our WRF-GS explicitly represents the propagation characteristics, as shown in Fig.
-2 (b).
-In addition, WRF-GS achieves higher reconstruction accuracy, higher sample efficiency, and faster rendering speed.
-Furthermore, we consider a downlink channel state information (CSI) prediction task in a multiple-input-multiple-output (MIMO) system as a case study to further demonstrate the effectiveness of WRF-GS in practical systems.
-Our results show that WRF-GS achieves better prediction performance of the downlink CSI than the baselines, outperforming NeRF2 by a significant margin of 2.43 dB.
-In summary, this paper makes the following contributions: i) We propose WRF-GS, a hybrid framework tailored for wireless channel modeling based on WRF reconstruction using neural networks and 3D-GS; ii) We adapt the optical 3D-GS to the RF domain by designing the modules of scenario representation network, projection model, and electromagnetic splatting.
-The integration of the EM wave physics and 3D-GS enables a visible, low-cost, and accurate channel modeling approach; iii) We conduct extensive experiments and a case study to demonstrate the effectiveness of WRF-GS.
-Numerical results show that WRF-GS outperforms existing methods.
+2 illustrates the synthesized spatial spectra of five algorithms based on the reconstructed WRF in a laboratory environment.
+Evidently, WRF-GS+ achieves the best prediction accuracy compared with the WRF-GS method and other baselines, closely matching the ground truth.
+We conduct extensive experiments to evaluate the WRF-GS and WRF-GS+ methods.
+The numerical results show that the proposed methods are capable of explicitly representing the propagation characteristics and achieving higher reconstruction accuracy, greater sample efficiency, and faster rendering speed in wireless channel modeling.
+Moreover, we present two case studies, i.e., the received signal strength indicator (RSSI) prediction and the downlink channel state information (CSI) prediction, to demonstrate the effectiveness of the proposed methods in practical systems.
+The RSSI prediction task focuses on estimating the signal strength at a specific location, while the CSI prediction task involves estimating the downlink CSI based on the uplink CSI measurements in a multiple-input-multiple-output (MIMO) system.
+The numerical results show that WRF-GS+ achieves superior performance in both tasks, outperforming existing methods by more than $0.7$ dB in RSSI prediction and $3.36$ dB in CSI prediction.
+Our main contributions are summarized as follows:
+  • We propose WRF-GS, a hybrid framework tailored for wireless channel modeling based on WRF reconstruction using neural networks and 3D-GS.
+This framework enables efficient and accurate modeling of wireless propagation characteristics.
+  • We adapt the optical 3D-GS to the RF domain by designing the modules of scenario representation network, projection model, and electromagnetic splatting.
+The integration of the EM wave physics and 3D-GS enables a visible, low-cost, and accurate channel modeling.
+  • We propose WRF-GS+, an enhanced version of WRF-GS, which employs deformable 3D Gaussians to separately model the static and dynamic components of the signals from virtual TXs.
+This approach improves efficiency and accuracy in characterizing the distribution of WRF.
+  • We conduct extensive experiments and two case studies to demonstrate the effectiveness of the WRF-GS and WRF-GS+ methods.
+Numerical results show that the proposed methods consistently outperform the existing methods.
+The remainder of this paper is organized as follows.
+Section II reviews the related works.
+Section III provides the preliminaries of the wireless channel model and 3D-GS.
+Sections IV and V detail the proposed WRF-GS and WRF-GS+ methods, respectively.
+Section VI presents the implementation and evaluation of the proposed methods in a laboratory environment using a real-world dataset.
+Two case studies are presented in Section VII to further demonstrate the practical applicability of the methods.
+Finally, Section VIII concludes the paper.
 II Related Works
@@
 Environment-aware channel modeling has attracted much attention since future wireless networks will generate abundant location-specific channel data and are expected to possess more powerful data mining and AI capabilities [14, 15].
-This motivated the adaptation of neural networks for more accurate channel modeling, i.e., learning the complex interactions between the environment and radio signals directly from location-based data [15].
+This motivated the adaptation of neural networks for more accurate channel modeling, i.e., learning the complex interactions between the environment and radio signals directly from location-based data [16, 17].
 To achieve this goal, a few recent works have been devoted to NeRF-based wireless channel modeling [18, 7, 8].
@@
 In this paper, we present WRF-GS for environment-aware channel modeling with 3D-GS, reconciling the accuracy and computational complexity.
+RSSI prediction is a typical task in wireless networks to estimate the signal strength at a specific location based on factors such as distance, environment, and prior measurements.
+This task is crucial for network management, indoor positioning, and signal coverage planning.
+Shin et al.
+[19] proposed an autonomous smartphone-based war-walking system (MRI) that automatically constructs a radio fingerprint database using limited user-collected data.
+MRI uses a log-distance path loss model to predict RSSI values throughout the entire indoor space based on partial measurements.
+Similarly, Zhao et al.
+[7] devised a turbo learning-enabled RSSI prediction method based on the NeRF2 framework using the Bluetooth Low Energy (BLE) dataset.
+In this paper, we apply the WRF-GS and WRF-GS+ methods to the RSSI prediction task by retraining them on the public BLE dataset [7].
 CSI prediction involves estimating future channel conditions based on current and historical data.
-Existing studies have been dedicated to predicting downlink channel states by observing uplink channels, as both are influenced by the same underlying physical environment and traverse identical paths [17, 18, 19].
+Existing studies have focused on predicting downlink channel states by observing uplink channels, as both are influenced by the same underlying physical environment and traverse identical paths [20, 21, 22].
 Liu et al.
@@
 [22] proposed R2F2 and OptML to predict the CSI using machine learning-based methods.
-In this paper, we apply WRF-GS to the CSI prediction task by transferring the estimated CSI from uplink to downlink.
+In this paper, we apply the WRF-GS and WRF-GS+ methods to the CSI prediction task by transferring the estimated CSI from uplink to downlink.
 III Preliminaries
-In this section, we present preliminaries on wireless channel modeling and 3D Gaussian Splatting.
+In this section, we present preliminaries on wireless channel modeling and 3D-GS technique.
 III-A Wireless Channel Modeling
-A generic wireless communication system typically consists of a TX that generates and modulates an information signal, which is then transmitted through the wireless channel to an RX.
-The transmitted signal can be represented as a complex number $s=Ae^{j\varphi}$, where $A$ and $\varphi$ denote the amplitude and phase of the transmitted signal, respectively.
-In the wireless medium, the propagation of EM signals is characterized by key phenomena such as path loss and multipath propagation.
-For the basic free space path loss model, the received signal $y$ at the RX is given by
+A generic wireless communication system typically consists of a TX that generates and modulates an information signal, which is then propagated through the wireless channel to an RX.
+The transmitted signal can be mathematically represented as a complex number $s=Ae^{j\varphi}$, where $A$ and $\varphi$ denote the amplitude and phase of the signal, respectively.
+In the wireless medium, the propagation of EM signals is influenced by fundamental physical phenomena, including path loss and multipath propagation.
+In the case of free space path loss, the received signal $y$ at the RX is expressed as:
 $y=Ae^{j\varphi}\cdot\Delta Ae^{j\Delta\varphi},$ (1)
-where $\Delta A$ and $\Delta\varphi$ are the amplitude attenuation factor and phase rotation experienced by the signal, respectively.
-Due to various propagation effects, including reflection, scattering, refraction, and diffraction, the received signal is the sum of multiple copies of the transmitted signal.
-Consequently, it is more generally expressed as
+where $\Delta A$ and $\Delta\varphi$ are the amplitude attenuation factor and phase rotation incurred during propagation, respectively.
+However, in realistic environments, the signal encounters a multitude of propagation effects, including reflection, scattering, refraction, and diffraction.
+These phenomena cause the transmitted signal to split into multiple replicas, each traversing different paths before arriving at the RX.
+As a result, the received signal is more accurately modeled as the superposition of these multipath components:
 $y=Ae^{j\varphi}\sum_{l=0}^{L-1}\Delta A_{l}e^{j\Delta\varphi_{l}},$ (2)
-where $L$ is the total number of propagation paths, and $\Delta A_{l}$ and $\Delta\varphi_{l}$ represent the amplitude attenuation factor and phase rotation associated with the $l$-th path, respectively.
-In multi-antenna systems, an antenna array is employed to characterize the energy distribution of the received signal $y$ from all directions, i.e., to estimate the AoA of the received signal.
+where $L$ denotes the total number of propagation paths, and $\Delta A_{l}$ and $\Delta\varphi_{l}$ correspond to the amplitude attenuation factor and phase rotation specific to the $l$-th path, respectively.
+This multipath formulation captures the complex interplay of signal interactions with the environment, providing a comprehensive framework to characterize the behaviors of wireless channels.
+In multi-antenna systems, an antenna array is employed to characterize the spatial energy distribution of the received signal $y$ from all directions, i.e., to estimate the AoA.
 To illustrate this concept, we consider an antenna array equipped with $\sqrt{K}\times\sqrt{K}$ antennas, as depicted in Fig.
 3.
-The spacing of two adjacent antennas is $D$, which is less than the wavelength $\lambda$.
+The spacing of two adjacent antennas is $D$, where $D<\lambda$ and $\lambda$ denotes the wavelength of the transmitted signal.
 As shown in Fig.
-3(b), the antenna array receives signals in a hemispherical plane to provide omnidirectional coverage and to efficiently capture signals from various directions.
-The direction of an RF source is characterized by the azimuthal angle $\alpha$ and elevation angle $\beta$, where $0^{\circ}\leq\alpha<360^{\circ}$ and $0^{\circ}\leq\beta<90^{\circ}$.
-The phase difference between the signals received by the $(m,n)$-th antenna pair is related to the distance of the two antennas from the source [20], i.e.,
+3(b), the antenna array operates within a hemispherical plane, providing omnidirectional coverage to efficiently capture signals from various directions.
+The direction of an RF source is characterized by two angles: the azimuthal angle $\alpha$ $(0^{\circ}\leq\alpha<360^{\circ})$ and the elevation angle $\beta$ $(0^{\circ}\leq\beta<90^{\circ})$.
+The phase difference between the signals received by the $(m,n)$-th antenna pair is determined by the relative distance of the two antennas from the source [23], expressed as:
 $\Delta\theta_{m,n}$ $=\mathrm{mod}\left(\frac{2\pi\Delta d_{m,n}}{\lambda},2\pi\right)$
 $=\mathrm{mod}\left(\frac{-2\pi r_{m,n}\cos\left(\alpha-\phi_{m,n}\right)\cos(\beta)}{\lambda},2\pi\right),$
-where $\Delta d_{m,n}$ is shown in Fig. 3(a) and can be computed by
+where $\Delta d_{m,n}$, illustrated in Fig. 3(a), is computed as
 $\Delta d_{m,n}=|BA_{m,n}|-|BA_{0,0}|=-r_{m,n}\cos(\alpha-\phi_{m,n})\cos\beta,$
-and $m,n=0,\ldots,\sqrt{K}-1$. In addition, $r_{m,n}=D\sqrt{m^{2}+n^{2}}$ and $\phi_{m,n}=\arctan 2(n,m)$ are the radius and angle of antenna $A_{m,n}$ at the polar coordinate system, respectively.
+Here, $m,n=0,\ldots,\sqrt{K}-1$, $r_{m,n}=D\sqrt{m^{2}+n^{2}}$ represents the radial distance of antenna $A_{m,n}$ in a polar coordinate system, and $\phi_{m,n}=\arctan 2(n,m)$ is its angular position.
 Figure 3: The AoA computation of the received signal with an antenna array.
@@
 (b) The spatial spectrum is generated from the antenna array.
-The illustration of spatial spectrum in Fig.
+The illustration of the spatial spectrum in Fig.
 2(c) is obtained by projecting the spatial spectrum to the X-Y plane.
-The AoA of the received signal is calculated by comparing the phases of signals received from multiple antennas.
-For AoA estimation, we use the antenna array to form a very narrow beam and steer it.
-When the beam is steered in the line-of-sight (LoS) direction, the array effectively blocks power from all other directions.
-This allows to reconstruct the spatial spectrum by steering the beam across a grid of azimuth and elevation angles and measuring the received signal power at each pointing direction.
-Specifically, we can form a matrix $\mathbf{P}$, where each element $\mathrm{P}(\alpha,\beta)$ represents the power of the received signal from the azimuthal angle $\alpha$ and elevation angle $\beta$, i.e.,
+The AoA of the received signal is calculated by comparing the phase differences of the signals received across multiple antennas.
+To achieve this, we use the antenna array to form a narrow beam and steer it across different directions.
+When the beam aligns with the line-of-sight (LoS) direction, the array effectively suppresses signals from other directions.
+This allows us to reconstruct the spatial spectrum by steering the beam across a grid of azimuth and elevation angles and measuring the received signal power at each pointing direction.
+Specifically, we can form a matrix $\mathbf{P}$, where each element $\mathrm{P}(\alpha,\beta)$ represents the received signal power from the azimuthal angle $\alpha$ and elevation angle $\beta$, i.e.,
 $\mathrm{P}(\alpha,\beta)=\left|\frac{1}{K}\sum_{m,n=0}^{\sqrt{K}-1,\sqrt{K}-1}e^{{j}\left(\Delta\hat{\theta}_{m,n}-\Delta\theta_{m,n}\right)}\right|^{2},\$ (3)
-and
+with
 $\Delta\hat{\theta}_{m,n}=\hat{\theta}_{m,n}-\hat{\theta}_{0,0},$
-where $\hat{\theta}_{m,n}$ denotes the measured phase of the RF signal received at the antenna $A_{m,n}$, and $\hat{\theta}_{0,0}$ is a constant term.
-Considering a one-degree angular resolution, the spatial spectrum can be reconstructed as a $360\times 90$ matrix:
+where $\hat{\theta}_{m,n}$ denotes the measured phase of the RF signal received at antenna $A_{m,n}$, and $\hat{\theta}_{0,0}$ is a constant reference phase.
+Considering a one-degree angular resolution, the spatial spectrum is reconstructed as a $360\times 90$ matrix:
 […]
-In summary, the reconstructed spatial spectrum can be interpreted as a spatial power distribution function, which directly corresponds to the WRF emitting from the transmitting source and propagating through the environment.
+The reconstructed spatial spectrum represents a spatial power distribution function, which directly corresponds to the WRF emitted by the transmitting source and propagated through the environment.
 In other words, the spatial spectrum serves as a representation of the underlying WRF, providing valuable insights into wireless channel modeling.
-Therefore, this paper aims to reconstruct the WRF and synthesize the spatial spectrum using the 3D-GS technique, which is introduced in the next subsection.
+Therefore, this paper aims to reconstruct the WRF and synthesize the spatial spectrum using the 3D-GS technique, which is introduced in the following subsection.
 III-B 3D Gaussian Splatting
-3D-GS is a technique that utilizes a set of 3D Gaussian functions to model the geometry and appearance of a scene, offering a compact and efficient representation [10].
-Unlike the computationally intensive and memory-heavy NeRF approach [5], 3D-GS provides a more lightweight and scalable solution for scene representation.
-Mathematically, 3D-GS provides a novel way to represent space using anisotropic Gaussian primitives, where each Gaussian is characterized by its own covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$ and center position vector $\boldsymbol{\mu}\in\mathbb{R}^{3}$.
-The 3D Gaussian distribution can be expressed as
+3D-GS is a technique that utilizes a set of 3D Gaussian functions to model the geometry and appearance of a scene, offering a compact and efficient representation [11, 24].
+Mathematically, 3D-GS represents space using anisotropic Gaussian primitives, where each Gaussian is characterized by its covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$ and center position vector $\boldsymbol{\mu}\in\mathbb{R}^{3}$.
+The 3D Gaussian distribution is expressed as
 $G(\boldsymbol{x})={e}^{-\frac{1}{2}(\boldsymbol{x}-\boldsymbol{\mu})^{T}\boldsymbol{\Sigma}^{-1}(\boldsymbol{x}-\boldsymbol{\mu})},$ (5)
-where $\boldsymbol{x}=(x_{0},x_{1},x_{2})$ represents the three-dimensional spatial position of the Gaussian point, and $\boldsymbol{\Sigma}$ can be expressed in terms of a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ as
+where $\boldsymbol{x}=(x_{0},x_{1},x_{2})$ represents the spatial coordinates of a Gaussian point in 3D space, and $\boldsymbol{\Sigma}$ can be expressed in terms of a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ as
 $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}.$ (6)
-The rendering process begins by transforming the 3D Gaussians $G(\boldsymbol{x})$ in space into 2D Gaussians $G^{\prime}(\boldsymbol{x^{\prime}})$ on the image plane [21].
-These 2D Gaussians are then sorted based on their depth information.
-Each tile is processed independently and in parallel for the 2D Gaussians in its coverage.
-Utilizing the opacity ${o}_{i}$ and color attribute $c_{i}$ of each Gaussian, alpha blending is applied to the color values on each pixel:
-$C=\displaystyle\sum_{i=1}^{N}{c}_{i}{\alpha}_{i}\displaystyle\prod_{j=1}^{i-1}(1-{\alpha}_{j})\ \mathrm{and}\ {\alpha}_{i}={o}_{i}{G}_{i}^{{}^{\prime}}(\varDelta\boldsymbol{p}_{i}),$ (7)
-where $N$ is the number of 3D Gaussians near a given pixel, and $\varDelta\boldsymbol{p}_{i}=\boldsymbol{p}_{i}-\boldsymbol{p}_{s}$ is the difference vector between the projected center $\boldsymbol{p}_{i}$ of a Gaussian and the sampling pixel position $\boldsymbol{p}_{s}$.
+The rendering process begins by projecting the 3D Gaussians $G(\boldsymbol{x})$ onto the image plane, resulting in 2D Gaussians $G^{\prime}(\boldsymbol{x^{\prime}})$ [25].
+These 2D Gaussians are then sorted based on their depth information to ensure proper occlusion handling.
+Each tile on the image plane is processed independently and in parallel for the 2D Gaussians within its coverage.
+Using the opacity ${o}_{i}$ and color attribute $c_{i}$ of each Gaussian, $\alpha$-blending is applied to compute the color values for each pixel:
+$C=\displaystyle\sum_{i=1}^{N}{c}_{i}{\alpha}_{i}\displaystyle\prod_{j=1}^{i-1}(1-{\alpha}_{j}),$ (7)
+where
+${\alpha}_{i}={o}_{i}{G}_{i}^{{}^{\prime}}(\varDelta\boldsymbol{p}_{i}).$
+Here, $N$ denotes the number of 3D Gaussians affecting a given pixel, and $\varDelta\boldsymbol{p}_{i}=\boldsymbol{p}_{i}-\boldsymbol{p}_{s}$ represents the difference vector between the projected center $\boldsymbol{p}_{i}$ of a Gaussian and the sampling pixel position $\boldsymbol{p}_{s}$.
 Once the color of each pixel is computed, the entire image is rendered.
-Subsequently, a comparison is made with the ground truth image to calculate the pixel-wise loss, which is then used to optimize the parameters of the model.
+The rendered image is then compared with the ground truth to calculate the pixel-wise loss, which is used to optimize the model parameters.
+In summary, 3D-GS is an efficient and scalable technique for representing and rendering 3D scenes in computer graphics by modeling scenes as a collection of Gaussian primitives.
+This motivates us to adopt the 3D-GS technique for wireless channel modeling based on explicit WRF reconstruction.
+This adaptation is achieved by treating Gaussian primitives as virtual TXs that capture wireless signal strengths and attenuation properties.
+By mapping the learned Gaussian representations onto the receiver’s hemispherical antenna perception plane, we can efficiently reconstruct the WRF and rapidly synthesize accurate spatial power spectra in Eqn.
+(4) and Fig.
+2(b).
 IV The WRF-GS Framework
-In this section, we present the WRF-GS framework for wireless channel modeling based on WRF reconstruction using 3D-GS.
-This novel framework harnesses the power of 3D-GS and adapts it from the optical domain to the RF domain.
-The key insight is to convert the space particles into virtual TXs using a continuous volumetric scene function in 3D-GS.
-IV-A Problem Description and Framework Overview
-We consider a scenario comprising a TX, an RX equipped with an antenna array, and various obstacles such as floors, walls, and furniture.
-The RX is placed at a fixed position, while the TX is randomly positioned in the space.
-Following [6], we assume that the distribution of main obstacles in the space is fixed, and any temporary perturbations caused by moving small obstacles are mitigated through techniques like Kalman filtering to minimize their impact.
-As described in Section III-A, the received signal is the superposition of sub-signals from multiple propagation paths.
-Each sub-signal can be represented as an LoS signal from a virtual TX.
-Our objective is to reconstruct the spatial power spectrum received at the RX antenna array by synthesizing the distribution of these virtual TXs within the scenario.
+In this section, we present the WRF-GS framework, a novel approach for wireless channel modeling based on the reconstruction of WRF using 3D-GS.
+This framework leverages the power of 3D-GS to explicitly characterize the spatial distribution of the WRF in a given environment.
+The key innovation lies in transforming spatial particles into virtual TXs using a continuous volumetric scene representation enabled by 3D-GS.
 Figure 4: An overview of the WRF-GS framework.
-The 3D point clouds, which can be randomly distributed or captured via LiDAR sensors, and the position of the TX are first passed into a scenario representation network.
+The 3D points, which can be randomly distributed or captured via LiDAR sensors, and the position of the TX are first passed into a scenario representation network.
 This network represents the virtual TXs in the scene using a set of 3D Gaussians, each of which carries environmental attenuation information and signal characteristics.
@@
 Finally, the resulting spatial spectra are rendered using the electromagnetic splatting method.
-This task is analogous to the optical 3D-GS technique that explicitly models the distribution of particles in a scene to enable arbitrary view synthesis [10].
-This motivates us to employ 3D-GS to reconstruct the WRF and propose a novel framework named WRF-GS.
-Specifically, WRF-GS aims to model the virtual TX distribution to synthesize the received spatial spectrum by combining the EM wave physics and 3D-GS.
-Since 3D Gaussians possess attributes such as volume, color, and opacity, and serve as the fundamental building blocks in the representation space, they can be naturally translated into virtual TXs in the RF domain.
-During EM propagation, the overall outcome is perceived as the combined effect of the signal and attenuation, as described in Eqn.
+IV-A Problem Description and Framework Overview
+We consider a wireless communication scenario comprising a TX, an RX equipped with an antenna array, and various obstacles such as floors, walls, and furniture.
+The RX is positioned at a fixed location, while the TX is randomly placed within the environment.
+Following [7], we assume that the primary obstacles in the environment remain stationary, and any transient perturbations caused by moving small obstacles are mitigated using techniques such as Kalman filtering to minimize their impact.
+As described in Section III-A, the received signal at the RX is a superposition of sub-signals from multiple propagation paths.
+Each sub-signal can be represented as a LoS signal originating from a virtual TX.
+Our goal is to reconstruct the spatial power spectrum received at the RX antenna array by modeling the distribution of these virtual TXs within the environment.
+This task is analogous to the optical 3D-GS technique, which explicitly models the distribution of particles in a scene to enable arbitrary view synthesis [11].
+This motivates us to adapt 3D-GS for WRF reconstruction and propose the novel WRF-GS framework.
+Specifically, WRF-GS models the distribution of virtual TXs to synthesize the received spatial spectrum by integrating the EM wave physics with 3D-GS.
+Since 3D Gaussians possess attributes such as volume, color, and opacity, serving as the fundamental building blocks in the representation space, they can be naturally reinterpreted as virtual TXs in the RF domain.
+During EM propagation, the overall signal is perceived as the combined effect of signal strength and attenuation, as described in Eqn.
 (2).
-In this context, the color attribute is replaced by the signal strength, while the opacity corresponds to the attenuation.
-Nevertheless, the projection and splatting of WRF reconstruction in 3D-GS are distinct from the optical domain described in Section III-B.
-In optical 3D-GS, the goal is to project 3D Gaussian particles representing objects onto the camera’s 2D image plane.
-This involves applying a coordinate projection transformation that accounts for the camera’s position, orientation, and intrinsic parameters.
-In contrast, the RF domain requires mapping the virtual TXs represented by 3D Gaussians onto a perception plane of the RX antenna array.
-Since the RX antenna array receives signals in a hemispherical plane, as discussed in Section III-A, a new projection approach is required to map the 3D Gaussians onto a hemispherical plane, rather than a flat 2D plane.
-This operation is crucial for aligning the virtual TXs with the specific direction of the RX antenna array.
-Furthermore, in optical 3D-GS, the rendered image is generated by the splatted Gaussians, including volume, color, and opacity.
-While the 3D Gaussian attributes in the RF domain represent the signal strength and attenuation.
-This necessitates a new splatting process to blend the contributions of all the virtual TXs, accounting for their signal strengths and attenuation, to synthesize the received spatial power spectrum.
-Our proposed WRF-GS effectively addresses the above challenges and enables the reconstruction of the spatial spectrum.
-The overall structure of WRF-GS is illustrated in Fig.
+In this context, the color attribute of a Gaussian corresponds to the signal strength, while the opacity represents the attenuation.
+However, the projection and splatting processes for WRF reconstruction in 3D-GS differ significantly from those in the optical domain, as described in Section III-B.
+In optical 3D-GS, the goal is to project 3D Gaussian particles representing objects onto a 2D image plane, which involves a coordinate projection transformation based on the camera’s position, orientation, and intrinsic parameters.
+In contrast, the RF domain requires mapping the virtual TXs represented by 3D Gaussians onto the perception plane of the RX antenna array.
+Since the RX antenna array receives signals in a hemispherical plane, as discussed in Section III-A, a new projection approach is needed to map 3D Gaussians onto a hemispherical plane rather than a flat 2D plane.
+This operation is critical for aligning the virtual TXs with the specific directional sensitivity of the RX antenna array.
+In addition, while optical 3D-GS renders images by blending the volume, color, and opacity of splatted Gaussians, the RF domain requires blending the signal strengths and attenuations of virtual TXs to synthesize the received spatial power spectrum.
+This necessitates a new splatting process tailored to the properties of EM wave propagation.
+Our proposed WRF-GS framework effectively addresses the aforementioned challenges and enables the accurate reconstruction of the spatial spectrum.
+The overall architecture of WRF-GS is illustrated in Fig.
 4.
-It takes the positions of physical TXs and 3D point clouds as input and outputs the spatial spectrum at the RX antenna array.
+It takes as input the positions of physical TXs and a set of randomly initialized 3D points representing spatial particles within the environment, and outputs the spatial spectrum at the RX antenna array.
 WRF-GS consists of the following three core modules:
-  • Scenario Representation Network: This module is responsible for representing the virtual TXs using 3D Gaussians.
-By processing the input data, the 3D Gaussians capture the signal strength and attenuation properties, which enables the synthesis of the WRF in the environment.
+  • Scenario Representation Network: This module represents the virtual TXs using 3D Gaussians.
+By processing the input data, the 3D Gaussians capture the signal strength and attenuation properties, enabling the synthesis of the WRF in the environment.
   • Projection Model: This model projects the virtual TXs represented by 3D Gaussians onto the perception plane of the RX antenna array.
-To account for the hemispherical nature of the antenna’s reception, we employ the Mercator projection to achieve this spatial projection.
-  • Electromagnetic Splatting: Based on the properties of 3D Gaussians and the characteristics of EM wave propagation in space, we employ a hardware acceleration algorithm to efficiently process and combine the contributions of the virtual TXs onto the RX antenna array, ultimately synthesizing the received spatial spectrum.
+To account for the hemispherical nature of the antenna’s reception, we employ the Mercator projection to achieve this spatial mapping.
+  • Electromagnetic Splatting: Leveraging the properties of 3D Gaussians and the characteristics of EM wave propagation, this module employs a hardware-accelerated algorithm to efficiently combine the contributions of the virtual TXs onto the RX antenna array, ultimately synthesizing the received spatial spectrum.
 In the following, we elaborate on these three modules.
 IV-B Scenario Representation Network
-The scenario representation network plays a crucial role in modeling the underlying relationships between the environment and radio signals by representing the virtual TXs as 3D Gaussians using two MLPs.
+The scenario representation network plays an important role in modeling the intricate relationships between the environment and radio signals by representing virtual TXs as 3D Gaussians using two MLPs.
 The network architecture, depicted in Fig.
-5, follows the DeepSDF structure [22].
-To capture the attenuation, which solely depends on the spatial location, we initially feed the positions of the initial point clouds (equivalent to the center coordinates of the 3D Gaussians $G(\boldsymbol{x})$) into a network consisting of eight fully connected layers.
-Each layer comprises ReLU activations and 128 channels.
-The output of this network provides the attenuation $\delta(\boldsymbol{x})$ of each input 3D Gaussain $G(\boldsymbol{x})$ at the corresponding position $\boldsymbol{x}$ and a feature vector.
-Subsequently, we pass this feature vector, along with the TX position $P_{\text{TX}}$, through another two-layer fully connected network (with ReLU activations and 128 and 64 channels, respectively).
-The final output is the signal $S(\boldsymbol{x})$ that represents the 3D Gaussian signal associated with the virtual TX at position $\boldsymbol{x}$.
-Therefore, for each 3D Gaussian $G(\boldsymbol{x})$ representing a virtual TX, we obtain the corresponding signal $S(\boldsymbol{x})$ and attenuation $\delta(\boldsymbol{x})$ from the scenario representation network, i.e.,
+5, is inspired by the DeepSDF structure [26] but tailored for wireless channel modeling.
+The network begins by processing the initial 3D points, which represent randomly sampled spatial particles within the environment.
+These points serve as the center coordinates of the 3D Gaussians $G(\boldsymbol{x})$ at location $\boldsymbol{x}$ and are used to initialize the Gaussian primitives.
+These positions are fed into a neural network to capture the attenuation properties, which depend solely on the spatial location.
+To characterize these properties, we introduce a new attenuation attribute $\delta(\boldsymbol{x})$ for each 3D Gaussian, which quantifies the signal attenuation caused by environmental interactions at position $\boldsymbol{x}$.
+This neural network consists of eight fully connected layers, each with ReLU activations and 128 channels.
+The output of this neural network includes the attenuation $\delta(\boldsymbol{x})$ of each input 3D Gaussain $G(\boldsymbol{x})$ and a feature vector that captures environmental information.
+Next, this feature vector, along with the TX position $P_{\text{TX}}$, is passed through a second two-layer fully connected network.
+The first layer has 128 channels and the second layer has 64 channels, both employing ReLU activations.
+The final output is the signal $S(\boldsymbol{x})$, which represents the 3D Gaussian signal associated with the virtual TX at position $\boldsymbol{x}$.
+Therefore, for each 3D Gaussian $G(\boldsymbol{x})$ representing a virtual TX, the scenario representation network produces the corresponding signal $S(\boldsymbol{x})$ and attenuation $\delta(\boldsymbol{x})$, as expressed by
 ${F}_{\Theta}:(G(\boldsymbol{x}),P_{\text{TX}})\Rightarrow\left(\delta(\boldsymbol{x}),S(\boldsymbol{x})\right),$ (8)
-where $\Theta$ denotes the learnable neural network weights.
-Considering that signals and attenuation in space are typically expressed as complex numbers, as shown in Eqn.
-(1), the two outputs here are denoted by $\delta(\boldsymbol{x})=\Delta A(\boldsymbol{x})e^{j\Delta\psi(\boldsymbol{x})}$ and $S(\boldsymbol{x})=A(\boldsymbol{x})e^{j\psi(\boldsymbol{x})}$, respectively, where $A(\boldsymbol{x})$ represents the amplitude and $\psi(\boldsymbol{x})$ represents the phase.
-This network is specifically designed to transform the 3D point clouds and TX position information into the signal and attenuation properties of 3D Gaussians, respectively.
-Unlike DeepSDF, which is typically applied to optics and assumes a fixed TX (light source) position, our network is designed to handle mobile TXs.
-Moreover, the outputs are all complex-valued signals that incorporate additional phase information.
-This unique feature enables our model to successfully represent both the signal characteristics of virtual TXs in space and the attenuation information of the surrounding environment.
+where $\Theta$ denotes the learnable weights of the network.
+Since signals and attenuation in space are typically expressed as complex numbers, as shown in Eqn.
+(1), the two outputs are expressed as $\delta(\boldsymbol{x})=\Delta A(\boldsymbol{x})e^{j\Delta\psi(\boldsymbol{x})}$ and $S(\boldsymbol{x})=A(\boldsymbol{x})e^{j\psi(\boldsymbol{x})}$, respectively, where $A(\boldsymbol{x})$ represents the amplitude and $\psi(\boldsymbol{x})$ represents the phase.
+This network is specifically designed to transform the 3D points (representing spatial particles) and TX position information into the signal and attenuation properties of 3D Gaussians.
+Unlike DeepSDF, which is typically applied in optic domains and assumes a fixed TX (light source) position, our network is uniquely adapted to handle mobile TXs.
+Moreover, the outputs are complex-valued signals that incorporate phase information, a critical feature for accurately representing wireless propagation.
+This capability enables our model to successfully capture both the signal characteristics of virtual TXs and the attenuation properties of the surrounding environment, making it well-suited for wireless channel modeling.
+IV-C Projection Model
+The projection model is responsible for mapping the virtual TXs represented by 3D Gaussians onto the perception plane of the RX antenna array.
+There are several key differences compared with the camera model used in optical 3D-GS.
+First, unlike pinhole or fisheye cameras, the RX antenna in our WRF-GS receives signals from a hemispherical direction.
+Second, we do not account for distortions caused by coordinate transformations, as we are only concerned with the signal at specific angular resolutions and ignore signals between discrete angles.
+Third, as shown in Eqn.
+(4), the arrangement of pixels in the final spatial spectrum differs from conventional photographs, which are designed for human visual perception.
+Instead, the spatial spectrum is structured to represent signal strength as a function of azimuth and elevation angles.
 Figure 5: Architecture of the neural model.
@@
 The second MLP then combines the spatial feature vector with the TX position to capture the signal characteristics.
-IV-C Projection Model
-The projection model maps the virtual TXs represented by 3D Gaussians onto the perception plane of the RX antenna array.
-There are several key differences compared with the camera model in optical 3D-GS.
-Unlike pinhole or fisheye cameras, the antenna in our WRF-GS receives signals from a hemispherical direction.
-In addition, we do not need to know the distortion caused by coordinate transformation.
-This is because we only care about the signal at the corresponding angular resolution and ignore the signal between angles.
-Moreover, as shown in Eqn.
-(4), the arrangement of pixels in the final spatial spectrum differs from conventional photographs that adhere to human visual perception.
-To project a 3D point $\boldsymbol{t}={[{t}_{x},{t}_{y},{t}_{z}]}^{T}$ onto the perception plane of the RX antenna array at $\boldsymbol{p}={[{p}_{x},{p}_{y}]}^{T}$, we employ the Mercator projection, a widely used orthographic cylindrical projection in cartography.
+To project a 3D point $\boldsymbol{t}={[{t}_{x},{t}_{y},{t}_{z}]}^{T}$ onto the perception plane of the RX antenna array at $\boldsymbol{p}={[{p}_{x},{p}_{y}]}^{T}$, we employ the Mercator projection, a widely used orthographic cylindrical projection in cartography [27].
 This allows us to map the latitude and longitude grid of a sphere onto a cylindrical surface, which can then be unrolled onto a flat plane.
-We define a right-handed Cartesian coordinate at the receiving end of the antenna, with the antenna array oriented in the positive Z-direction and the signal being received in a spherical plane, as shown in Fig.
+We define a right-handed Cartesian coordinate system at the receiving end of the antenna, with the antenna array oriented in the positive Z-direction and signals received in a spherical plane, as shown in Fig.
 LABEL:Proj(a).
-Using longitude $\Omega_{\text{lon}}$ and latitude $\Omega_{\text{lat}}$, we can concisely and accurately represent positional relationships in space, which is expressed using inverse trigonometric functions as:
+Using longitude $\Omega_{\text{lon}}$ and latitude $\Omega_{\text{lat}}$, we can concisely and accurately represent positional relationships in space.
+By applying inverse trigonometric functions, we obtain
 $\begin{bmatrix}\Omega_{\text{lon}}\\ \Omega_{\text{lat}}\end{bmatrix}=\begin{bmatrix}\arctan 2({t}_{y}/{t}_{x})\\ \arcsin({t}_{z}/{t}_{r})\end{bmatrix},$ (9)
-where ${t}_{r}=\sqrt{{t}_{x}^{2}+{t}_{y}^{2}+{t}_{z}^{2}}$ denotes the distance from the center coordinate of the 3D Gaussian in space to the center of the antenna coordinate, $\arctan 2(\cdot)$ is the 4-quadrant inverse tangent.
-Since we only consider signals from the upper hemisphere, we have $-\pi\leq\Omega_{\text{lon}}<\pi$ and $20\leq\Omega_{\text{lat}}<\pi/$ 2, and the lower hemisphere is ignored in the following coordinate transformations.
-Subsequently, we convert the latitude-longitude coordinates (see Fig.
+where ${t}_{r}=\sqrt{{t}_{x}^{2}+{t}_{y}^{2}+{t}_{z}^{2}}$ denotes the distance from the center coordinate of the 3D Gaussian in space to the origin of the antenna coordinate system.
+In addition, the function $\arctan 2(\cdot)$ is the 4-quadrant inverse tangent.
+Since we only consider signals from the upper hemisphere, the longitude and latitude ranges are constrained to $-\pi\leq\Omega_{\text{lon}}<\pi$ and $20\leq\Omega_{\text{lat}}<\pi/$ 2.
+Signals from the lower hemisphere are ignored in subsequent coordinate transformations.
+Next, we convert the latitude-longitude coordinates (see Fig.
 LABEL:Proj(b)) into uniform coordinates (see Fig.
-LABEL:Proj(c)) as
+LABEL:Proj(c)) as follows:
 $\begin{bmatrix}{s}_{x}\\ {s}_{y}\end{bmatrix}=\begin{bmatrix}\Omega_{\text{lon}}/\pi\\ 2\Omega_{\text{lat}}/\pi\end{bmatrix},$ (10)
-such that $-1\leq{s}_{x}<1$ and $10\leq{s}_{y}<$ 1. The final step of the projection is to transform the uniform coordinates into pixel coordinates of arbitrary resolution (see Fig. LABEL:Proj(d)) as
+such that $-1\leq{s}_{x}<1$ and $10\leq{s}_{y}<$ 1.
+The final step of the projection is to transform the uniform coordinates into pixel coordinates of arbitrary resolution (see Fig.
+LABEL:Proj(d)), i.e.,
 $\begin{bmatrix}{p}_{x}\\ {p}_{y}\end{bmatrix}=\begin{bmatrix}({s}_{x}+1)\times W/2\\ {s}_{y}\times H\end{bmatrix},$ (11)
-where $W$ and $H$ represent the number of pixels contained in the width and height of the image, respectively. Considering a one-degree angular resolution, we set $W=360$ and $H=90$.
+where $W$ and $H$ represent the number of pixels contained in the width and height of the image, respectively. For a one-degree angular resolution, we set $W=360$ and $H=90$.
 Through the above coordinate transformations, we maintain the relationship between the signal angles, ensuring that each pixel value in the spatial spectrum corresponds to the AoA of the signal.
-This allows us to calculate the signals arriving in each direction at the RX antenna array in the next subsection.
+This allows us to calculate the signal strength arriving from each direction at the RX antenna array, as detailed in the following subsection.
 IV-D Electromagnetic Splatting
-After the coordinate projection transformation, the 3D Gaussians representing the virtual TXs are mapped onto the 2D plane.
+After the coordinate projection transformation, the 3D Gaussians representing the virtual TXs are mapped onto a 2D plane.
 Note that each 2D Gaussian may cover multiple pixels.
-In other words, the value of each pixel is the result of the combined contributions from multiple 2D Gaussians.
-Therefore, the role of electromagnetic splatting module is to efficiently compute the value of each pixel in the 2D plane, which is then utilized to reconstruct the spatial spectrum.
+In other words, the value of each pixel is determined by the combined contributions of multiple overlapping 2D Gaussians.
+The role of the electromagnetic splatting module is to efficiently compute the value of each pixel on the 2D plane, which is then used to reconstruct the spatial spectrum.
 As illustrated in Fig.
-9, electromagnetic splatting incorporates the EM wave propagation into the differentiable rasterisation algorithm [10].
-To facilitate fast parallel computing, we group adjacent pixels into non-overlapping “tiles” that can be computed independently, as depicted in Fig.
+9, electromagnetic splatting incorporates the principles of EM wave propagation into a differentiable rasterisation algorithm [11].
+To enable fast parallel computing, we group adjacent pixels into non-overlapping “tiles”, which are processed independently, as depicted in Fig.
 9(a).
-For each 2D Gaussian, we first determine which tiles it covers, duplicate the Gaussian, and assign the copies to the corresponding tiles.
-These 2D Gaussians are then sorted by depth order within each tile, as shown in Fig.
+For each 2D Gaussian, we first identify the tiles it covers, duplicate the Gaussian, and assign these copies to the corresponding tiles.
+The 2D Gaussians within each tile are then sorted by depth, as shown in Fig.
 9(b).
 In the scenario representation network, we obtain the signal $S(\boldsymbol{x}_{i})$ and signal attenuation $\delta(\boldsymbol{x}_{i})$ of each 3D Gaussian $G(\boldsymbol{x}_{i})$ at an arbitrary position $\boldsymbol{x}_{i}$.
-Assuming there are a total of $N$ 3D Gaussians contributing to a particular angle, we can express the effect of the $i^{th}$ Gaussian $G(\boldsymbol{x}_{i})$ on the RX by combining Eqn.
+Assuming there are a total of $N$ 3D Gaussians contributing to a specific angle (pixel), we can express the effect of the $i^{th}$ Gaussian $G(\boldsymbol{x}_{i})$ on the RX by combining Eqn.
 (2) and Eqn.
@@
 $S_{i}(\boldsymbol{x})=\left(\prod_{j=0}^{i-1}\delta(\boldsymbol{x}_{j})\right)S(\boldsymbol{x}_{i}).$ (12)
-Consequently, for all 3D Gaussians associated with an angle (pixel) $k$ , the combined effect is expressed as
+Consequently, for all 3D Gaussians associated with a specific angle (pixel) $k$ , the combined effect is expressed as
 $R_{k}=\sum_{i=1}^{N}S_{i}(\boldsymbol{x}).$ (13)
@@
 For the pixel value $R_{2}$ in the upper-right corner, the contributions from $G(\boldsymbol{x}_{1})$, $G(\boldsymbol{x}_{2})$, and $G(\boldsymbol{x}_{3})$ must be considered.
-The signal from $G(\boldsymbol{x}_{1})$ has no attenuation, but the signals from $G(\boldsymbol{x}_{2})$ and $G(\boldsymbol{x}_{3})$ are attenuated by $\delta(\boldsymbol{x}_{1})$ and $\delta(\boldsymbol{x}_{1})\delta(\boldsymbol{x}_{2})$, respectively, due to the depth ordering.
-The computation for the remaining pixel values $R_{k}$ follows a similar principle.
+The signal from $G(\boldsymbol{x}_{1})$ experiences no attenuation, while the signals from $G(\boldsymbol{x}_{2})$ and $G(\boldsymbol{x}_{3})$ are attenuated by $\delta(\boldsymbol{x}_{1})$ and $\delta(\boldsymbol{x}_{1})\delta(\boldsymbol{x}_{2})$, respectively, due to the depth ordering.
+The computation for the remaining pixel values $R_{k}$ follows the same principle.
+In simple terms, we consider each projected Gaussian as a virtual TX that can affect multiple angles, while signals from multiple virtual TXs are received at each angle.
+Depending on their distance to the RX, these signals undergo varying degrees of attenuation, ultimately determining the received signal strength at each angle.
+As explained in Section III-A, the reconstructed spatial spectrum can be represented as a spatial power distribution function, where the value is proportional to the square of the amplitude of the received signal.
+By deriving the received signal for each angle using Eqn.
+(12) and Eqn.
+(13), we obtain the spatial spectrum described in Eqn.
+(4), which serves as a representation of the underlying WRF.
+IV-E Summary
+The WRF-GS framework integrates physical models and deep learning techniques to enable comprehensive environment-aware channel modeling and high-fidelity spatial spectrum synthesis.
+At the core of this approach lies the establishment of a connection between virtual TXs and 3D Gaussians.
+The scenario representation network models the spatial distribution of signal propagation and attenuation using 3D Gaussians based on input data.
+To transform these 3D spatial representations onto a 2D plane, we employ the Mercator map projection, ensuring accurate spatial mapping.
+Finally, the electromagnetic splatting module employs a parallel and efficient algorithm to compute pixel values for each direction, enabling rapid synthesis of the spatial spectrum.
+V The WRF-GS+ Framework
+The proposed WRF-GS framework demonstrates remarkable effectiveness in reconstructing the WRF and synthesizing spatial spectra for wireless channel modeling, as shown in Section VI and Section VII.
+However, WRF-GS faces limitations in accurately modeling high-frequency signal variations, as the static nature of the 3D Gaussians in WRF-GS restricts its ability to capture rapid changes in the channel caused by complex multipath effects.
+To address this challenge, we propose WRF-GS+, an enhanced framework that introduces deformable 3D Gaussians to better model dynamic signal variations and incorporates electromagnetic splatting with $\alpha$-blending to simplify the modeling process.
+V-A Scenario Representation Network Using Deformable 3D Gaussians
+While WRF-GS effectively models wireless environments, it struggles to accurately capture the rapid signal variations caused by complex multipath effects.
+To address this limitation, we introduce deformable 3D Gaussians, which enables the framework to decompose the signal into static components and dynamic components.
+This decomposition allows the network to separately model large-scale fading (e.g., path loss) and small-scale fading (e.g., multipath effects), significantly improving its ability to handle high-frequency signal changes.
+As illustrated in Fig.
+10, the scenario representation network initializes Gaussian primitives from 3D points, each assigned with attributes such as opacity, signal strength, rotation, and scaling.
+Unlike WRF-GS, which introduces an additional attribute to explicitly model signal attenuation, WRF-GS+ directly leverages the opacity attribute to capture attenuation characteristics.
+The signal strength, rotation, and scaling attributes are determined exclusively by the center position of the 3D Gaussian.
+Notably, these static components of the signal, corresponding to large-scale fading, depend solely on the relative position of the virtual TX to the RX, as the environmental distribution remains invariant.
+These attributes are thus used to characterize the static components: signal strength is represented by the original color attribute, while rotation and scaling define the size and orientation of Gaussians, thereby affecting the coverage area of the virtual TXs.
 Figure 9: An illustration of electromagnetic splatting.
@@
 (13).
-In simple terms, we consider each projected Gaussian as a virtual TX that can affect multiple angles, while signals from multiple virtual TXs are received at each angle.
-Depending on the distance of these signals to the RX, they will be attenuated to different degrees, ultimately resulting in the received signal on each angle.
-As explained in Section III-A, the reconstructed spatial spectrum can be represented as a spatial power distribution function, where the value is proportional to the square of the amplitude of the received signal.
-Therefore, by deriving the received signal on each angle from Eqn.
+The dynamic components, which correspond to small-scale fading, are modeled using a deformation network ${D}_{\Theta}$.
+This network captures the variations in signal propagation paths caused by complex multipath effects, leading to changes in the coverage and shape of virtual TXs.
+The deformation network, implemented as an MLP, maps the coordinates of the 3D Gaussians and the TX position to their corresponding deviations in signal strength, rotation, and scaling.
+The weights $\Theta$ of the MLP are optimized through this mapping process.
+As shown in Fig.
+11, the deformation network processes the input through eight fully connected layers with ReLU activations and 256-dimensional hidden layers, producing a 256-dimensional feature vector.
+This vector is then passed through three additional fully connected layers (without activation functions) to separately output the offsets for signal strength, rotation, and scaling.
+Notably, similar to NeRF, we concatenate the feature vector output from the fourth layer with the input and feed them into the fifth layer to enhance feature representation.
+Given the TX position, it is fed into the network along with the 3D Gaussians to generate the offsets:
+${D}_{\Theta}:(G(\boldsymbol{x}),P_{\text{TX}})\Rightarrow\left(\Delta_{\text{sig}}(\boldsymbol{x}),\Delta_{\text{rot}}(\boldsymbol{x}),\Delta_{\text{scal}}(\boldsymbol{x})\right),$ (14)
+where $\Delta_{\text{sig}}(\boldsymbol{x})$, $\Delta_{\text{rot}}(\boldsymbol{x})$, and $\Delta_{\text{scal}}(\boldsymbol{x})$ represent the offsets in signal strength, rotation, and scaling, respectively.
+Once the dynamic components are obtained, they are combined with the static components to form the complete properties of the 3D Gaussians at a specific TX position.
+These properties are then used in the subsequent coordinate projection transformation and electromagnetic signal calculation processes.
+By incorporating deformable 3D Gaussians, the improved scenario representation network effectively integrates the characteristics of large-scale and small-scale fading into the channel model.
+This enhancement enables the network to accurately capture the high-frequency variations, significantly improving its interpretability and characterization capabilities.
+Figure 10: The enhanced scenario representation network.
+The random 3D points are used to initialize 3D Gaussians, each characterized by properties such as position, opacity, signal strength, rotation, and scale.
+The 3D Gaussians, along with the TX position, are fed into the deformation network to compute dynamic offsets for the properties.
+These dynamic offsets are combined with the original static properties to derive the final characteristics of 3D Gaussians.
+These updated properties are then processed through the projection model and electromagnetic splatting to synthesize the final spatial spectra.
+V-B Electromagnetic Splatting with $\alpha$-blending
+With the introduction of a new structure in the scenario representation network, WRF-GS+ enhances the original framework by associating environmental information with the opacity attribute of 3D Gaussians.
+This opacity is translated into the weight of each virtual TX, which implicitly characterizes attenuation information, enabling a more accurate representation of signal propagation.
+Compared with WRF-GS, WRF-GS+ eliminates the need for additional attributes.
+This simplifies the 3D-GS model by reducing the number of parameters and memory overhead while improving convergence.
+In the field of optics, opacity is modeled as particle density, which characterizes the probability of particles in space being blocked.
+This concept is adapted to the wireless domain, where opacity is used to represent the environmental influence on signal propagation.
+Specifically, the opacity of a 3D Gaussian reflects the material properties at its location, which in turn determines the weight of the corresponding virtual TX.
+By incorporating this property, WRF-GS+ can effectively model how environmental factors, such as obstacles or reflectors, affect signal strength at different locations.
+Based on the $\alpha$-blending technique described in Eqn.
+(7) for optical rendering, we extend this concept to the electromagnetic domain by adapting it to complex number calculations.
+This allows us to rewrite Eqn.
 (12) and Eqn.
-(13), we obtain the spatial spectrum described in Eqn.
-(4), which serves as a representation of the underlying WRF.
-IV-E Summary
-The WRF-GS framework integrates physical and deep learning models to enable comprehensive environment-aware channel modeling and high-fidelity spatial spectrum synthesis.
-At the core of this approach lies the establishment of a connection between virtual TXs and 3D Gaussians.
-The scenario representation network models the spatial distribution of signal propagation and attenuation using 3D Gaussians based on input data.
-To transform these 3D spatial representations onto a 2D plane for further processing, we employ the Mercator map projection, ensuring accurate spatial mapping.
-Finally, the electromagnetic splatting module employs a parallel and efficient algorithm to compute pixel values in each direction, thereby achieving fast spatial spectrum synthesis.
-V WRF-GS Implementation and Evaluation
-In this section, we implement and evaluate WRF-GS in a laboratory environment using a real-world dataset in [6].
-V-A Implementation
-WRF-GS is implemented in a Python environment using modified custom CUDA kernels for rasterization [10] and is trained on an NVIDIA GeForce RTX 3090 GPU.
-Detailed settings and implementation are provided below.
-Position Initialization: We use the LiDAR 3D point clouds [^1] Note that WRF-GS also works well by randomly initializing the point clouds within a specified range of space.
-shown in Fig.
-2(a) to initialize the positions of 3D Gaussian points.
-The reason is that the most significant signal attenuation usually occurs on the object’s surface [7], which corresponds to the 3D point clouds.
-This initialization enables a fast convergence of WRF-GS.
-Besides, we adopt an adaptive density control strategy [10] to learn the position distribution of the 3D Gaussians during the training process.
+(13) as:
+$R_{k}=\displaystyle\sum_{i=1}^{N}({S}(\boldsymbol{x}_{i})+\Delta_{\text{sig}}(\boldsymbol{x}_{i})){\alpha}_{i}\displaystyle\prod_{j=1}^{i-1}(1-{\alpha}_{j}),$ (15)
+where $R_{k}$ represents the received signal at a specific angle (pixel) $k$.
+This formula indicates that the received signal is the superposition of contributions from Gaussians, weighted by their respective $\alpha$-blending coefficients.
+Each Gaussian’s contribution consists of a static component ${S}(\boldsymbol{x}_{i})$ and a dynamic component $\Delta_{\text{Sig}}(\boldsymbol{x}_{i})$, reflecting the influence of environmental and TX movement effects, respectively.
+To provide an intuitive understanding of this process, we revisit the example illustrated in Fig.
+9(c).
+For the pixel value $R_{1}$ in the upper-left corner, only the signal from $G(\boldsymbol{x}_{1})$ contributes, but it is now weighted by its opacity coefficient $\alpha_{1}$.
+For the pixel value $R_{2}$ in the upper-right corner, the contributions from $G(\boldsymbol{x}_{1})$, $G(\boldsymbol{x}_{2})$, and $G(\boldsymbol{x}_{3})$ must be considered.
+The coefficient for the signal from $G(\boldsymbol{x}_{1})$ remains $\alpha_{1}$, while the coefficients for $G(\boldsymbol{x}_{2})$ and $G(\boldsymbol{x}_{3})$ are adjusted to $\alpha_{2}(1-\alpha_{1})$ and $\alpha_{3}(1-\alpha_{1})(1-\alpha_{2})$, respectively.
+The calculation for other pixel values $R_{k}$ follows the same principle.
+By leveraging the static nature of environmental information, we associate this information with the opacity attribute of 3D Gaussians, which depends only on their positions.
+WRF-GS+ achieves a more accurate and interpretable representation of the WRF by decomposing the virtual TX into static and dynamic components.
+Figure 11: The architecture of the deformation network, designed to learn the dynamic components in the scene.
+The 3D Gaussians and the TX position are first encoded using positional encoding and then fed into the deformation network.
+The network outputs the offsets in signal strength, rotation, and scaling attributes for each 3D Gaussian, enabling the representation of dynamic signal variations caused by TX movement and environmental interactions.
+V-C Summary
+The proposed WRF-GS+ framework addresses the limitations of WRF-GS in modeling dynamic signal variations and parameter redundancy.
+First, leveraging insights from the different characteristics of large-scale fading and small-scale fading in wireless communication scenarios, we enhance the scenario representation network by introducing deformable 3D Gaussians.
+This approach can efficiently capture the high-frequency signal variations caused by complex multipath effects.
+By integrating deep learning with physics-based modeling, the network achieves a more effective representation of high-frequency signal variations.
+Second, we enhance the electromagnetic splatting process by using the inherent opacity of 3D Gaussians as the weighting factor, eliminating the need for the network to explicitly represent attenuation parameters.
+This modification not only reduces computational complexity but also enhances the framework’s efficiency by leveraging the natural properties of 3D Gaussians to model signal attenuation implicitly.
+Finally, we implement an $\alpha$-blending method to compute the received signals from each direction in parallel, synthesizing the spatial spectrum with improved efficiency.
+VI Implementation and Evaluation
+In this section, we implement and evaluate the WRF-GS and WRF-GS+ methods in a laboratory environment using a real-world dataset in [7].
+VI-A Implementation Details
+The proposed methods are implemented in Python, utilizing customized CUDA kernels for rasterization [11], and trained on an NVIDIA GeForce RTX 3090 GPU.
+It is worth noting that we handle complex signals by separating them into real and imaginary components via Euler’s formula rather than decomposing complex signals into amplitude and phase.
+This operation enables parallelized computations in the CUDA kernels.
+The implementation details are provided below.
+Position Initialization: We randomly generate 3D points within a specified range.
+The positions of these points are used to initialize the center coordinates of the 3D Gaussians.
+In addition, we implement an adaptive density control strategy [11] to learn the positional distribution of the 3D Gaussians during the training process.
+Compared with the LiDAR point-cloud initialization [1], the random point initialization can overcome the requirement for multimodal input and minimize the reliance on the quality of LiDAR point clouds.
 Position Encoding: To improve the spatial resolution, we introduce an efficient position encoding method [6], i.e.,
 $\gamma(\boldsymbol{t})=(\sin(\pi\boldsymbol{t}),\cos(\pi\boldsymbol{t}),\ldots,\sin(2^{L}\pi\boldsymbol{t}),\cos(2^{L}\pi\boldsymbol{t})),$ (16)
-where $\boldsymbol{t}$ represents the 3D coordinates of the TX and $L$ is the order of the position encoding.
+where $\boldsymbol{t}$ represents the 3D coordinates input and $L$ is the order of the position encoding.
 For the input position encoding of $P_{\text{TX}}$ and $G(\boldsymbol{x})$ in the MLPs, we set $L=9$.
-Optimization Details: In the scenario representation network, we need to train two MLPs and learn the 3D Gaussian representations using the training dataset.
-We adopt the Adam optimizer for network training and 3D Gaussian representation.
-The loss function is defined as the difference between the synthesized spatial spectrum $I_{\text{pred}}$ and the ground truth $I_{\text{gt}}$, i.e.,
+Optimization Details: In the scenario representation network, we need to train the MLPs and learn the 3D Gaussian representations using the training dataset.
+We adopt the Adam optimizer for training and 3D Gaussian representation.
+The loss function is computed by comparing the difference between the synthesized spatial spectrum $I_{\text{pred}}$ and the ground truth $I_{\text{gt}}$:
 $\mathcal{L}=(1-\eta)|I_{\text{gt}}-I_{\text{pred}}|+\eta(1-\xi(I_{\text{gt}},I_{\text{pred}})),$ (17)
-where the $\xi(I_{\text{gt}},I_{\text{pred}})$ is the structural similarity index measure (SSIM) function [23], which is used to measure the similarity between two images.
+where $\xi(I_{\text{gt}},I_{\text{pred}})$ is the structural similarity index measure (SSIM) function [28], which is used to measure the similarity between two images.
 In addition, $\eta=0.2$ is a weighting factor.
-V-B Experiment Design
+VI-B Experiment Design
 Dataset.
-We use an open-source dataset provided in NeRF2 [6] to evaluate the WRF-GS framework.
-This dataset was obtained in a laboratory environment, which consists of multiple tables, shelves, and several small rooms.
+We use an open-source dataset provided in NeRF2 [7] to evaluate the proposed frameworks.
+This dataset is obtained in a laboratory environment, which consists of multiple tables, shelves, and several small rooms.
 The 3D LiDAR point clouds of the environment are shown in Fig.
@@
 To collect the training dataset, the position of the RX is fixed and the position of the TX is systematically varied throughout the laboratory.
-Each sample in the dataset consists of the TX position and the corresponding spatial spectrum measured at the RX (i.e., Eqn (4)).
-Benchmarks. We compare WRF-GS with several baselines:
+Each sample in the dataset consists of the TX position and the corresponding spatial spectrum measured at the RX (i.e., Eqn.
+(4)).
+Benchmarks. We compare WRF-GS+ and WRF-GS with several baselines:
   • Ray Tracing[3]: This approach emits rays from the RX through every pixel in the laboratory using the 3D LiDAR point clouds.
-These rays interact with objects to simulate the propagation of signal in the real world.
+These rays interact with the objects, resulting in reflection, diffraction, and scattering, which are recursively traced to simulate real-world ray propagation using the RayTracing toolkit in MATLAB.
   • Variational Autoencoder (VAE) [29]: VAE is commonly used to generate new data samples by mapping observed data to a continuous latent space [30].
@@
 The generator learns to generate realistic samples to deceive the discriminator, while the discriminator aims to distinguish the real samples from the generated ones.
-  • NeRF2 [6]: NeRF2 is a method for spatial spectrum synthesis using the NeRF technique. After training, NeRF2 can predict the spatial spectrum at arbitrary TX positions.
-V-C Evaluation
-We first evaluate the accuracy of the spatial spectrum synthesis generated by the WRF-GS, NeRF2, VAE, and ray tracing methods, as shown in Fig.
-2(c).
+  • NeRF2 [7]: NeRF2 is a method for spatial spectrum synthesis using the NeRF technique.
+Similarly to our work, NeRF2 can synthesize the spatial spectrum at arbitrary TX positions after training in a given scenario.
+VI-C Evaluation
+We first compare the accuracy of the spatial spectrum synthesis generated by the WRF-GS+, WRF-GS, NeRF2, VAE, and ray tracing methods, as shown in Fig.
+2(b).
 The visualization is obtained by transforming the matrix form in Eqn.
 (4) into a polar coordinate representation.
-The spatial spectrum of WRF-GS for the case when the TX is located at position P3 is a projection of the constructed WRF visualized in Fig.
-2(b).
-The results in Fig.
-2(c) show that WRF-GS achieves the best prediction accuracy, closely matching the ground truth.
-Fig.
-10 compares the SSIM [23] values of the synthesized spatial spectra generated by different methods.
-SSIM values range from 0 to 1, with higher values indicating greater similarity between the synthesized and groud truth samples.
+From Fig.
+2(b) we see that WRF-GS+ achieves the best performance compared with the WRF-GS method and the baselines, closely matching the ground truth.
+Next, we compare the SSIMs of the synthesized spatial spectra generated by different methods.
+The value of SSIM ranges from 00 to $1$, where a higher value indicates greater similarity between the synthesized and ground truth samples.
 The cumulative distribution function (CDF) of SSIM illustrates the disparities between the methods and the ground truth.
 From Fig.
-10, we see that WRF-GS outperforms the baselines.
-Specifically, the median SSIM values for WRF-GS, NeRF2, VAE, DCGAN, and ray tracing are 0.82, 0.78, 0.70, 0.56, and 0.38, respectively.
-The slightly inferior performance of NeRF2 is due to the challenge in enumerating the virtual TX locations along the propagation paths for integration or rendering.
-In contrast, the ray tracing method achieves the worst performance, as it heavily relies on accurate environmental material information.
-Figure 10: The SSIM values of the synthesized spatial spectra generated by the WRF-GS, NeRF2, VAE, DCGAN, and ray tracing methods.
-Figure 11: The median SSIM values of the WRF-GS and NeRF2 methods under different training dataset sizes.
+12, we see that the median SSIM values for WRF-GS+, WRF-GS, NeRF2, VAE, DCGAN, and ray tracing are $0.90$, $0.82$, $0.78$, $0.70$, $0.56$, and $0.38$, respectively, and their 90th percentiles are $0.95$, $0.88$, $0.86$, $0.82$, $0.69$, and $0.61$.
+Evidently, WRF-GS+ significantly outperforms WRF-GS thanks to its adaptation of deformable 3D Gaussians in the scenario representation network design.
+This operation enables the WRF-GS+ method to efficiently characterize the distribution of the WRF in a given scene by separately modeling the static and dynamic components of the signals from virtual TXs.
+In addition, the poor performance of NeRF2 is due to its reliance on the voxel sampling rate along the pixel ray.
+As a result, a high resolution results in significant computational overhead and demands extensive training data, while a lower resolution compromises the representation of WRF.
+Thus, there is a trade-off between representation capability and computational efficiency.
+In contrast, the ray tracing method achieves the poorest performance as it heavily relies on environmental material information, which is often hard to obtain.
+Figure 12: The SSIM values of the synthesized spatial spectra generated by the WRF-GS+, WRF-GS, NeRF2, VAE, DCGAN, and ray tracing methods.
+Figure 13: The median SSIM values of the WRF-GS+, WRF-GS and NeRF2 methods under different training dataset sizes.
 Fig.
-11 presents the median SSIM values of the WRF-GS and NeRF2 methods across varying sizes of the training dataset.
+13 compares the median SSIM values of the WRF-GS+, WRF-GS, and NeRF2 methods across varying sizes of the training dataset.
 We maintain a consistent test dataset and randomly subsample the original training dataset to create new training datasets ranging from $200$ to $4,800$ samples.
-The results demonstrate that the WRF-GS framework consistently outperforms the NeRF2 method for all training dataset sizes, improving the median SSIM by $4\%$ to $8\%$.
-More importantly, the performance gap increases as the training dataset size decreases.
-This suggests that WRF-GS is well-suited for practical applications in channel modeling, as it requires less training data and, consequently, less time and effort from experienced engineers.
-Figure 12: The rendering time required to synthesize a spatial spectrum using different methods under the same settings.
+It can be seen that WRF-GS+ consistently outperforms the WRF-GS and NeRF2 methods for all training dataset sizes, improving the median SSIM by $3\%$ to $13\%$ and $11\%$ to $19\%$, respectively.
+This demonstrates that WRF-GS+ has the highest sample efficiency.
+More importantly, the performance gap increases as the training dataset size decreases, which indicates that the proposed methods require less training data and, consequently, less time and effort from experienced engineers.
+Figure 14: The rendering time required to synthesize a spatial spectrum using different methods.
 Fig.
 14 illustrates the rendering time required to synthesize a spatial spectrum using different methods under the same settings.
-The results show that, for processing a single sample, WRF-GS takes $0.005$s, while NeRF2, VAE, DCGAN, and Ray Tracing takes $0.2$s, $0.05$s, $0.09$s, and $7$s, respectively.
-The processing time of WRF-GS is usually one to three orders of magnitude less than other methods.
-This superior computational efficiency makes WRF-GS well-suited for latency-sensitive applications.
-VI Case Study: CSI Prediction
-In this section, we demonstrate the effectiveness of WRF-GS on the downlink CSI prediction task in MIMO systems.
+The results show that WRF-GS+ achieves a processing time of $0.008$s per sample, while WRF-GS, NeRF2, VAE, DCGAN, and Ray Tracing require $0.005$s, $0.2$s, $0.05$s, $0.09$s, and $7$s, respectively.
+Notably, the processing times of WRF-GS+ and WRF-GS are one to three orders of magnitude lower than those of the other methods.
+This significant advantage makes the proposed methods highly suitable for latency-sensitive applications.
+Although the processing time of WRF-GS+ is slightly higher than that of WRF-GS, this difference is primarily due to the inclusion of the deformation network in WRF-GS+, which introduces more parameters compared to the two MLPs used in WRF-GS.
+However, given the significant improvement in spectrum synthesis accuracy achieved by WRF-GS+, this millisecond-level difference is negligible and acceptable in most practical application scenarios.
+VII Case Studies
+In this section, we further demonstrate the effectiveness of the proposed methods on RSSI and CSI prediction tasks.
+VII-A Case Study I: RSSI Prediction
+We first apply the WRF-GS and WRF-GS+ methods to the RSSI prediction task, which involves estimating the signal strength at a specific location based on factors such as distance, environmental conditions, and prior measurements.
+RSSI prediction is essential for network management, indoor positioning, and signal coverage planning.
+For example, the RSSI prediction accuracy directly affects the positioning accuracy of the localization task and the performance of the base station (BS) deployment in cellular networks.
+Unlike the previous spatial spectrum reconstruction task, the RSSI measurements are obtained by a single omnidirectional antenna instead of an antenna array.
+As a result, the output is a composed signal containing the power from all directions rather than the $90\times 360$ pixel matrix.
+To complete the RSSI prediction task, the proposed methods are trained on the public BLE dataset [7].
+We compare the proposed methods with the NeRF2 method and the MRI [19] method.
+MRI is a learning-based method that interpolates the RSSI values at the unsampled location using a log-distance path-loss model.
+VII-A1 Experiment Setup
+The BLE dataset is obtained in a scene containing several rooms with a total area of about $15,000ft^{2}$.
+There are $21$ RXs and a TX moving in this area.
+The RSSI at these RXs is recorded when the TX moves from one position to another.
+There are a total of $6,000$ RSSI measurements in the BLE dataset.
+Note that we consider the measurement to be an invalid value when the RX is far from the TX.
+In this situation, the recorded RSSI is set to $-100$ dBm.
+There are $2,000$ valid RSSI measurements for each RX, where $1,600$ measurements are used for training and the remaining $400$ are used for testing.
+VII-A2 RSSI Prediction Accuracy
+Figure 15: The RSSI prediction errors of the WRF-GS+, WRF-GS, NeRF2, and MRI methods on the public BLE dataset.
+Fig.
+15 compares the RSSI prediction errors of the WRF-GS+, WRF-GS, NeRF2, and MRI methods on the public BLE dataset.
+The prediction error is defined as the absolute difference between the predicted and ground truth RSSI values.
+In simulation, we train the proposed methods on the dataset of each RX.
+The result measured in dB is the average value of the RSSI prediction errors for $12$ RXs.
+From Fig.
+15, we see that WRF-GS+ has the lowest RSSI prediction error among the WRF-GS, NeRF2, and MRI methods.
+The median of WRF-GS+ is $2.4$ dB ($10$-th percentile: $1.1$ dB; $90$-th percentile: $4.6$ dB).
+In contrast, the median errors of WRF-GS, NeRF2, and MRI are $2.9$ dB, $3.1$ dB, and $8.3$ dB, respectively.
+The proposed methods clearly outperform the baselines, indicating that the 3D-GS-based method is more suitable than the NeRF-based method in this task.
+The superior performance of the WRF-GS+ is due to its ability to separate the signals from virtual TXs into static and dynamic components.
+This operation enables an accurate RSSI prediction even with continuous TX movements.
+VII-B Case Study II: CSI Prediction
+We adapt the WRF-GS+ and WRF-GS methods to the downlink CSI prediction task in MIMO systems.
 Typically, the downlink and uplink transmission are operated on different frequency bands in frequency domain duplex systems, where the uplink-downlink reciprocity does not hold.
-The base station (BS) can obtain the downlink CSI of each antenna from the uplink channel feedback by sending pilot sequences to the end devices.
+The BS can obtain the downlink CSI of each antenna from the uplink channel feedback by sending pilot sequences to the end devices.
 However, the overhead of such feedback is linearly proportional to the number of antennas and devices, making it impossible in massive MIMO systems.
@@
 As pointed out in [32], each CSI is unique and highly correlated with the physical environment.
-Therefore, there is a mapping between the device’s location and its uplink CSI.
+Therefore, there is a mapping between the device’s location and its uplink CSI, which is similar to fingerprint-based localization.
 Accounting for this, Eqn.
@@
 where $I_{u}(\boldsymbol{x})$ is the uplink CSI at location $\boldsymbol{x}$.
-Thus, we can retrain a WRF-GS to perform the downlink CSI prediction task.
-In this case, the input is changed from the TX’s location to the uplink CSI, and the output is changed from the spatial spectrum to the downlink CSI.
-VI-A Experiment Setup
-We adopt the public Argos channel dataset [28] to train the MLPs and 3D Gaussian representations in WRF-GS for the downlink CSI prediction task.
-This dataset is collected in a real environment and the CSI is measured in different environments where a BS is equipped with $104$ antennas and serves multiple users.
+Thus, we can apply the WRF-GS framework to the downlink CSI prediction task by employing a similar configuration in Section VI-A.
+For the WRF-GS+ method, we adapt the deformation network to this CSI prediction task by modifying Eqn. (14) into
+${D}_{\Theta}:(G(\boldsymbol{x}),I_{u}(\boldsymbol{x}))\Rightarrow\left(\delta_{\text{Sig}}(\boldsymbol{x}),\delta_{\text{r}}(\boldsymbol{x}),\delta_{\text{s}}(\boldsymbol{x})\right).$ (19)
+Note that the input is changed from the TX’s location to the uplink CSI, and the output of the whole pipeline is changed from the spatial spectrum to the downlink CSI.
+Figure 16: Comparison of the ground truth CSI and the predicted CSI using the proposed methods in terms of amplitude and phase via different sub-carriers.
+VII-B1 Experiment Setup
+We train the WRF-GS and WRF-GS+ methods on the public Argos channel dataset [33] for the downlink CSI prediction task.
+This dataset is collected in a real environment.
+The CSI is measured in different environments where a BS is equipped with $104$ antennas and serves multiple users.
 Each CSI measurement contains $52$ subcarriers.
-Similar to previous works [6, 17], we regard the first $26$ subcarriers as the uplink channel and the remaining $26$ subcarriers as the downlink channel.
-Apart from NeRF2 [6], we also compare WRF-GS with the following three approaches:
+Similarly to [7] and [20], we regard the first $26$ subcarriers as the uplink channel and the remaining $26$ subcarriers as the downlink channel.
+Unlike the spatial spectrum reconstruction task, which requires computing power from all directions separately, we focus on a complex value of the predicted CSI at each subcarrier.
+Notice that this complex value is a composed signal that sums in all directions.
+Apart from NeRF2, we also compare the proposed methods with the following baselines:
   • FIRE [20]: FIRE is a deep learning method based on the VAE architecture and is used to predict the downlink CSI by learning the latent distribution of the uplink CSI.
@@
   • OptML [22]: OptML is a deep learning method based on the R2F2 framework and can be used to predict the underlying path information of the downlink channel.
-VI-B Channel Estimation Accuracy
+VII-B2 Channel Estimation Accuracy
 Fig.
-13 depicts the prediction accuracy of the WRF-GS framework for the downlink CSI prediction task.
-The blue solid line represents the input ground truth uplink (GT Uplink) CSI, the orange solid line represents the ground truth downlink (GT Downlink) CSI, and the black dashed line is the predicted (Pred Downlink) CSI output by WRF-GS.
-It is seen that the predicted result is very close to the ground truth.
-Figure 13: The comparison of the ground truth CSI and the predicted CSI using WRF-GS in terms of the channel amplitude and phase via different sub-carries.
-To quantify the prediction accuracy, we use the SNR metric:
-$\text{SNR}=-10\log_{10}\left({\left\|\mathrm{}C_{\text{pred}}-C_{\text{GT}}\right\|^{2}}/{\left\|C_{\text{GT}}\right\|^{2}}\right),$ (17)
+16 depicts the prediction accuracy of the WRF-GS and WRF-GS+ methods in the downlink CSI prediction task.
+Since the predicted CSI is a complex value, we measure the prediction accuracy on the amplitude and phase parts.
+We see that the predicted results are close to the ground truth.
+This means that the proposed methods can successfully predict the downlink CSI by extracting information from the uplink CSI based on the neural network architecture and 3D-GS.
+Figure 17: The prediction accuracy of different methods in the downlink CSI prediction task.
+The figure shows the maximum, minimum, median, upper and lower quartile ranges of the CEA distribution for different methods, as well as the outliers.
+The higher the CEA, the better the prediction accuracy.
+Next, we adopt the following metric to quantify the prediction accuracy and refer to it as CEA, representing the channel estimation accuracy in dB:
+$\text{CEA}=-10\log_{10}\left(\frac{\left\|\mathrm{}C_{\text{pred}}-C_{\text{GT}}\right\|^{2}}{\left\|C_{\text{GT}}\right\|^{2}}\right),$ (20)
 where $C_{\text{pred}}$ is predicted downlink channel and $C_{\text{GT}}$ is ground truth downlink channel.
-The higher the SNR, the better the prediction accuracy.
+The higher the CEA, the better the prediction accuracy.
 Fig.
-14 shows the SNRs of five CSI prediction algorithms.
-Notably, WRF-GS achieves a median SNR of $22.98$ dB ($10$th percentile: $20.82$ dB, $90$th percentile: $24.33$ dB).
-The median SNRs of NeRF2, FIRE, R2F2, and OptML are $20.55$ dB, $15.29$ dB, $8.57$ dB, and $8.47$ dB, respectively.
-This demonstrates that WRF-GS improves the CSI prediction accuracy by efficiently modeling wireless channels.
-Figure 14: The prediction accuracy of different methods in the downlink CSI prediction task.
-The figure shows the maximum, minimum, median, upper and lower quartile ranges of the SNR distribution for different methods, as well as the outliers.
-The higher the SNR, the better the prediction accuracy.
-VII Conclusion
-This paper introduced WRF-GS, a novel framework that combines 3D-GS and neural networks to enable fast and accurate wireless channel modeling.
-To adapt the 3D-GS technique from the optical domain to the RF domain, we designed three key modules for WRF-GS: a scenario representation network, a projection model, and an electromagnetic splatting technique.
-These modules allow WRF-GS to accurately capture the underlying physics of EM wave propagation.
-Remarkably, WRF-GS can synthesize new channel characteristics within milliseconds using a small number of training samples.
-Experimental results showed that WRF-GS outperforms existing methods in spatial spectrum reconstruction and excels in the CSI prediction task, achieving an improvement of over 2.43 dB.
-We envision that WRF-GS will enable various applications in future networks by leveraging its high-fidelity visualization, low sample complexity, and rapid rendering speed.
-The authors have provided public access to their code at https://github.com/wenchaozheng/WRF-GS.
+17 compares the prediction accuracy of the proposed methods and baselines.
+The results show that WRF-GS+ achieves a median CEA of $23.91$ dB ($10$th percentile: $20.85$ dB, $90$th percentile: $25.52$ dB), while WRF-GS achieves a median CAE of $22.98$ dB ($10$th percentile: $20.82$ dB, $90$th percentile: $24.33$ dB).
+In contrast, the median CEAs of NeRF2, FIRE, R2F2, and OptML are $20.55$ dB, $15.29$ dB, $8.57$ dB, and $8.47$ dB, respectively.
+The proposed methods consistently outperform the baselines.
+Notably, WRF-GS+ achieves the highest prediction accuracy, surpassing the best baseline NeRF2 by a significant margin of $3.36$ dB.
+This demonstrates that the superior capability of the proposed methods in efficiently and accurately modeling wireless channels.
+VIII Conclusion
+This paper first introduced WRF-GS, a novel framework for wireless channel modeling based on WRF reconstruction using neural network and 3D-GS.
+WRF-GS effectively captures the interactions between the environment and radio signals by refining the optical 3D-GS technique from three key aspects: scenario representation network, projection model, and electromagnetic splatting.
+To improve efficiency and accuracy, we further proposed WRF-GS+ by incorporating deformable 3D Gaussians and $\alpha$-blending to the scenario representation network and electromagnetic splatting modules, respectively.
+Experimental results showed that the proposed methods outperform the baselines for spatial spectrum synthesis.
+They can synthesize new channel characteristics within milliseconds using a small number of training samples.
+Moreover, WRF-GS+ achieves superior performance in both RSSI and CSI prediction tasks, surpassing existing methods by more than $0.7$ dB and $3.36$ dB, respectively.
+We envision that the proposed methods will enable various applications in future networks by leveraging their high-fidelity visualization, low sample complexity, and rapid rendering speed.